### v1

# Epidemic threshold and localization of the SIS model on directed complex networks

Abstract
We study the susceptible-infected-susceptible (SIS) model on directed complex networks within the quenched mean-field approximation. Combining results from random matrix theory with an analytic approach to the distribution of fixed-point infection probabilities, we derive the phase diagram and show that the model exhibits a nonequilibrium phase transition between the absorbing and endemic phases for $c\geq\lambda^{-1}$, where $c$ is the mean degree and $\lambda$ the average infection rate. Interestingly, the critical line is independent of the degree distribution but is highly sensitive to the form of the infection-rate distribution. We further show that the inverse participation ratio of infection probabilities diverges near the epidemic threshold, indicating that the disease may become localized on a small fraction of nodes. These results provide a comprehensive picture of how network heterogeneities shape epidemic spreading on directed contact networks.
preprint: APS/123-QED
I Introduction
The susceptible-infected-susceptible (SIS) model on networks provides a minimal yet powerful framework to investigate the interplay between epidemic spreading and the structure of the underlying contact network Kiss et al. (2017). In the SIS model, each node (individual) can be in either a susceptible or infected state. A susceptible node $i$ can become infected by a neighbour $j$ at a rate $\lambda_{ij}$, typically assumed to be uniform, $\lambda_{ij}=\lambda$. An infected node, in turn, becomes healthy at a rate conventionally set to unity. The average fraction of infected individuals, referred to as the prevalence, is the natural order-parameter for characterizing the spread of the epidemic. In the limit of an infinitely large number $N$ of individuals, the SIS model may exhibit a nonequilibrium phase transition between an absorbing phase Pastor-Satorras et al. (2015), in which the epidemic dies out, and an endemic phase, characterized by a stationary state with a nonzero prevalence.
A central problem in the study of epidemic spreading is to understand how the network structure influences the epidemic threshold $\lambda_{\rm c}$ Chakrabarti et al. (2008); Chatterjee and Durrett (2009); Castellano and Pastor-Satorras (2010); Ferreira et al. (2012); Goltsev et al. (2012); Li et al. (2013); Van Mieghem and van de Bovenkamp (2013); Kwon and Kim (2013), which separates the absorbing phase ($\lambda\leq\lambda_{\rm c}$) from the endemic phase ($\lambda>\lambda_{\rm c}$). The quenched mean-field (QMF) theory Ferreira et al. (2012); Mata and Ferreira (2013); Pastor-Satorras et al. (2015) provides an effective approximation for the SIS model by neglecting dynamical correlations between neighboring nodes Mata and Ferreira (2013), leading to a set of coupled dynamical equations for the single-node infection probabilities. Within this framework, the epidemic threshold for a network of size $N$ is given by Ferreira et al. (2012); Van Mieghem and van de Bovenkamp (2013)
$\lambda_{\rm c}(N)=1/\Lambda_{1}(N),$ (1)
where $\Lambda_{1}(N)>0$ denotes the leading eigenvalue (spectral radius) of the network adjacency matrix Newman (2010). Beyond its role in QMF theory, Eq. (1) provides a rigorous lower bound for the exact epidemic threshold on any given network Van Mieghem and van de Bovenkamp (2013).
The spectral properties of the adjacency matrix thus play a fundamental role in the dynamics of the SIS model. For undirected networks, the epidemic threshold follows from well-established analytic results for the leading eigenvalue $\Lambda_{1}(N)$ Krivelevich and Sudakov (2003); Chung et al. (2003). In such networks, the expected value of $\Lambda_{1}(N)$ typically scales with the maximum degree. Consequently, for degree distributions with unbounded support, the epidemic threshold vanishes in the limit $N\rightarrow\infty$ Castellano and Pastor-Satorras (2010); Ferreira et al. (2012). This prediction of the QMF theory is fully consistent with rigorous results for the SIS model Chatterjee and Durrett (2009), which prove the absence of an absorbing phase for any nonzero infection rate in undirected networks.
Another spectral property relevant to the dynamics of the SIS model is the spatial localization of the leading eigenvector associated with $\Lambda_{1}(N)$ Goltsev et al. (2012); Pastor-Satorras and Castellano (2018). A strongly localized eigenvector has a finite number of nonzero components as $N\rightarrow\infty$ Abou-Chacra et al. (1973); Mirlin (2000), whereas an extended eigenvector is characterized by an extensive number of nonzero components. Localization effects are expected to be particularly relevant near the epidemic threshold, where the leading eigenvector closely approximates the stationary endemic state, providing insights into network-based immunization strategies that target influential nodes Pastor-Satorras and Castellano (2018). In undirected networks, the localization of the leading eigenvector has been studied through the inverse participation ratio (IPR) Goltsev et al. (2012); Pastor-Satorras and Castellano (2016, 2018); Liu and Van Mieghem (2019), defined in terms of the fourth moment of the eigenvector components. Numerical results for the IPR indicate that, in finite-size undirected networks, the leading eigenvector is localized on a vanishing fraction of nodes Goltsev et al. (2012); Pastor-Satorras and Castellano (2016, 2018). In particular, for degree distributions that decay sufficiently fast, this localization occurs at the hub with the largest degree Goltsev et al. (2012); Pastor-Satorras and Castellano (2016).
In contrast to undirected networks, the SIS model on directed contact networks, where the infection rates $\lambda_{ij}$ are unidirectional  [^1] In a directed contact network, if node $j$ infects node $i$ with rate $\lambda_{ij}$, the opposite rate $\lambda_{ji}$ is zero., remains poorly understood. Spreading models on directed networks provide a natural framework for studying the transmission of computer viruses through email networks Newman et al. (2002); Ebel et al. (2002), the diffusion of information in social networks Krueger et al. (2024), and the transmission of diseases between patients and health care workers Meyers et al. (2006). Only a few works have examined how directionality impacts the SIS model Li et al. (2013); Kwon and Kim (2013). Numerical studies on finite networks with both directed and bidirected edges have shown that $\lambda_{\rm c}(N)$ increases with the fraction of directed links Li et al. (2013). In addition, the SIS model has been studied on directed networks using heterogeneous mean-field theory Kwon and Kim (2013), which does not fully capture the underlying network structure. As a result, even basic problems, such as how the directed network structure affects the epidemic threshold, remain unresolved.
In this work, we study the QMF theory of the SIS model on directed networks with arbitrary distributions of degrees and infection rates. We determine the epidemic threshold $\lambda_{\rm c}$ in the limit $N\rightarrow\infty$, showing that this model undergoes an absorbing phase transition as a function of the network structure. Our calculation of $\lambda_{\rm c}$ partially relies on random-matrix analytic results for the leading eigenpair of directed weighted networks Metz and Pérez Castillo (2016); Lucas Metz et al. (2019); Metz and Neri (2021). In contrast with undirected networks, the leading eigenvalue of directed networks remains finite as $N\rightarrow\infty$ Metz and Pérez Castillo (2016), even when the degree distribution has unbounded support.
We show that the epidemic threshold is determined by the leading eigenvalue only when the network parameters are such that the gap between the leading and the subleading eigenvalue remains finite as $N\rightarrow\infty$. With the main goal of computing $\lambda_{\rm c}$ for arbitrary combinations of network parameters, we derive an exact equation for the full distribution of the stationary infection probabilities in the limit $N\rightarrow\infty$, using the cavity method from spin-glass theory Mézard and Parisi (2001); Metz et al. (2010). The numerical solutions of this equation yield both the epidemic threshold and the IPR as functions of the network structure. We find that $\lambda_{\rm c}$ is independent of the degree distribution, whereas the shape of the infection-rate distribution strongly influences the epidemic threshold, in particular for large fluctuations in the infection rates. Furthermore, we show that the IPR diverges near the epidemic threshold as $N\rightarrow\infty$, in agreement with analytic predictions derived from the moments of the leading eigenvector Metz and Neri (2021). This divergence arises from a large fraction of nodes with infection probabilities that strongly fluctuate near zero, indicating that the disease becomes localized on a vanishing fraction of network nodes.
The paper is organized as follows. In the next section, we introduce the SIS model on directed complex networks within the QMF approximation. In section III, we derive the equation for the stationary distribution of infection probabilities using the cavity method. The results for the epidemic threshold and the IPR as functions of the network parameters are discussed in section IV. Finally, we present a summary of our findings and concluding remarks in section V.
II The SIS model on directed networks
We consider the SIS model in the quenched mean-field approximation Pastor-Satorras et al. (2015); Mata and Ferreira (2013). The probability $\rho_{i}(t)$ ($i=1,\dots,N$) that a node $i$ is infected at time $t$ evolves according to
$\frac{d\rho_{i}}{dt}=-\rho_{i}(t)+\left[1-\rho_{i}(t)\right]\sum_{j=1(\neq i)}^{N}A_{ij}\rho_{j}(t),$ (2)
where the elements $\{A_{ij}\}_{i,j=1,\dots,N}$ of the $N\times N$ weighted adjacency matrix $\bm{A}$ have the form $A_{ij}=C_{ij}\lambda_{ij}$, with $C_{ij}\in\{0,1\}$ and $\lambda_{ij}>0$. The binary variables $\{C_{ij}\}_{i,j=1,\dots,N}$ specify the structure of the contact network: if $C_{ij}=1$, there is a directed link $j\rightarrow i$ from node $j$ to $i$, whereas $C_{ij}=0$ otherwise. The indegree $K_{i}$ (outdegree $L_{i}$) of node $i$ is given by $K_{i}=\sum_{j=1}^{N}C_{ij}$ ($L_{i}=\sum_{j=1}^{N}C_{ji}$). We consider directed random networks generated from the configuration model Newman et al. (2001); Fosdick et al. (2018), where the degree sequences $K_{1},\dots,K_{N}$ and $L_{1},\dots,L_{N}$ are independent and identically distributed random variables drawn from $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$, with $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ denoting, respectively, the indegree and outdegree distributions. The average degree $c$ is determined from
$c=\sum_{k=0}^{\infty}k\,p_{{\rm in},k}=\sum_{\ell=0}^{\infty}\ell\,p_{{\rm out},\ell}.$ (3)
We will present results for three examples of degree distributions: Poisson, geometric and power-law Newman et al. (2001).
The coefficient $\lambda_{ij}>0$ denotes the infection rate that node $j$ infects $i$. We assume that $\{\lambda_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables sampled from a distribution $P_{\lambda}(x=\lambda_{ij})$ with mean $\lambda$ and variance $\sigma^{2}$. We will present results for two choices of $P_{\lambda}(x)$. The first one is the $\Gamma$-distribution
$P_{\lambda,{\rm g}}(x)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-\beta x}\Theta(x),$ (4)
with parameters
$\alpha=\lambda^{2}/\sigma^{2}\quad\text{and}\quad\beta=\lambda/\sigma^{2},$ (5)
and with $\Theta(x)$ representing the Heaviside step function. The second example is the Pareto distribution with finite variance,
$P_{\lambda,{\rm p}}(x)=\frac{\gamma x_{0}^{\gamma}}{x^{\gamma+1}}\Theta\left(x-x_{0}\right),$ (6)
where
$\gamma=1+\sqrt{1+\frac{\lambda^{2}}{\sigma^{2}}}>2\quad\text{and}\quad x_{0}=\left(\frac{\sqrt{1+\frac{\lambda^{2}}{\sigma^{2}}}}{1+\sqrt{1+\frac{\lambda^{2}}{\sigma^{2}}}}\right)\lambda.$ (7)
Note that both distributions are solely parametrized in terms of $\lambda$ and $\sigma$. For $\sigma=0$, the infection rates are uniform ($\lambda_{ij}=\lambda$).
The network ensemble is fully specified by the distributions $p_{k\ell}$ and $P_{\lambda}(x)$. For $c>1$, the network contains a giant strongly connected component Kryven (2016), ensuring that the spectrum of $\bm{A}$ has a continuous component in the limit $N\rightarrow\infty$ Neri and Metz (2020). This is the interesting regime where the nodes strongly interact with each other and the disease can eventually infect a finite fraction of individuals.
Our goal is to understand how the structure of the contact network influences the stationary states of the model. Setting $\frac{d\rho_{i}}{dt}=0$ in Eq. (2), we obtain the fixed-point equations
$\rho_{i}=\frac{\sum_{j=1}^{N}A_{ij}\rho_{j}}{1+\sum_{j=1}^{N}A_{ij}\rho_{j}}\quad(i=1,\dots,N).$ (8)
To characterize the stationary behaviour and the phase diagram in the limit $N\rightarrow\infty$, we study the distribution
$\mathcal{P}(\rho)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\delta(\rho-\rho_{i})$ (9)
of the fixed-point infection probabilities $\rho_{1},\dots,\rho_{N}$. The moments of $\mathcal{P}(\rho)$ read
$\langle\rho^{n}\rangle=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\rho_{i}^{n}=\int_{0}^{1}d\rho\mathcal{P}(\rho)\rho^{n}.$ (10)
In particular, the prevalence $\langle\rho\rangle$ quantifies the average fraction of infected individuals.
In the next section, we develop an analytic approach, based on the cavity method of spin-glass theory Mézard and Parisi (2001); Metz et al. (2010), that yields the following self-consistent equation for the distribution $\mathcal{P}(\rho)$,
$\mathcal{P}(\rho)=\frac{1}{Z}\sum_{k=0}^{\infty}p_{{\rm in},k}\left(\prod_{j=1}^{k}\int_{0}^{1}d\rho_{j}\mathcal{P}(\rho_{j})\int_{0}^{\infty}dx_{j}P_{\lambda}(x_{j})\right)$
$\times\delta\Bigg{[}(1-\rho)\sum_{j=1}^{k}x_{j}\rho_{j}-\rho\Bigg{]},$ (11)
where the constant $Z$ ensures that $\int_{0}^{1}d\rho\mathcal{P}(\rho)=1$. This equation can be also inferred from the effective dynamics obtained in Metz (2025) using dynamical mean-field theory.
Equation (11) can be numerically solved through a Monte-Carlo iterative method known as population dynamics Mézard and Parisi (2001); Mézard and Montanari (2009); Metz and Pérez Castillo (2016); Lucas Metz et al. (2019). The method is based on the discretization of $\mathcal{P}(\rho)$ in terms of a population with $M$ stochastic variables $\rho_{1},\dots,\rho_{M}$. In the standard version of the algorithm Mézard and Parisi (2001); Mézard and Montanari (2009); Metz and Pérez Castillo (2016), each iteration proceeds by randomly selecting a member of the population and updating it according to the constraint imposed by the Dirac-$\delta$ in Eq. (11). After a sufficient number of iterations, the population $\rho_{1},\dots,\rho_{M}$ converges to a stationary profile, yielding an approximate solution of Eq. (11). In the present context, the core of the algorithm remains unchanged, but we have to include the weight arising from the normalization factor $Z$, which determines the number of elements updated in parallel at each iteration. The details of this modified version of the algorithm are discussed in Lucas Metz et al. (2019).
III The cavity approach for the distribution of infection probabilities
In this section, we present a first-principles derivation of Eq. (11) for the distribution of the fixed-point infection probabilities $\rho_{1},\dots,\rho_{N}$ in the limit $N\rightarrow\infty$. This section can be skipped by readers primarily interested in the results.
The variables $\bm{\rho}=(\rho_{1},\dots,\rho_{N})^{T}$ solve the equations
$F_{i}(\bm{\rho})=-\rho_{i}+(1-\rho_{i})\sum_{j=1}^{N}A_{ij}\rho_{j}=0,$ (12)
where $i=1,\dots,N$. Assuming that, for a single realization of $\bm{A}$, Eq. (12) admits a unique solution, the joint probability density of $\bm{\rho}$ can be formally written as Kurchan (1991)
$\mathcal{P}_{N}(\bm{\rho})=\dfrac{\prod_{i=1}^{N}\delta\left[F_{i}(\bm{\rho})\right]}{\int\limits_{0}^{1}\left(\prod_{i=1}^{N}d\rho_{i}\right)\prod_{i=1}^{N}\delta\left[F_{i}(\bm{\rho})\right]}.$ (13)
Using the Fourier transform of the Dirac-$\delta$, we write $\mathcal{P}_{N}(\bm{\rho})$ as
[…]
where
$\gamma_{N}(\bm{\rho},\hat{\bm{\rho}})=\frac{\exp{\left[H_{N}(\bm{\rho},\hat{\bm{\rho}})\right]}}{\int\limits_{-\infty}^{\infty}\left(\prod_{i=1}^{N}d\hat{\rho}_{i}\right)\int\limits_{0}^{1}\left(\prod_{i=1}^{N}d\rho_{i}\right)\exp{\left[H_{N}(\bm{\rho},\hat{\bm{\rho}})\right]}}$, (15)
with
$H_{N}(\bm{\rho},\hat{\bm{\rho}})=-i\sum_{j=1}^{N}\hat{\rho}_{j}\rho_{j}+i\sum_{lj=1}^{N}A_{lj}\hat{\rho}_{l}(1-\rho_{l})\rho_{j}.$ (16)
Our purpose is to compute the local marginals $\{\mathcal{P}_{N,i}(\rho_{i})\}_{i=1}^{N}$ on the network nodes by using the cavity method Metz et al. (2010).
From Eq. (14), the marginal $\mathcal{P}_{N,i}(\rho_{i})$ is given by
$\mathcal{P}_{N,i}(\rho_{i})=\int\limits_{-\infty}^{\infty}d\hat{\rho}_{i}\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}).$ (17)
Let us extract the $i$-th term from the sums in Eq. (16), namely
$H_{N}(\bm{\rho},\hat{\bm{\rho}})=-i\hat{\rho}_{i}\rho_{i}+i\rho_{i}\sum_{j\in\partial_{i}^{\rm out}}\lambda_{ji}\hat{\rho}_{j}(1-\rho_{j})$
[…]
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}\int\limits_{-\infty}^{\infty}\Bigg{(}\prod_{j\in\partial_{i}}d\hat{\rho}_{j}\Bigg{)}\int\limits_{0}^{1}\Bigg{(}\prod_{j\in\partial_{i}}d\rho_{j}\Bigg{)}$
$\times\exp{\Bigg{[}i\rho_{i}\sum_{j\in\partial_{i}^{\rm out}}\lambda_{ji}\hat{\rho}_{j}(1-\rho_{j})+i\hat{\rho}_{i}(1-\rho_{i})\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}\Bigg{]}}$
$\times\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}),$ (19)
where $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}})$ is defined on the neighbourhood $\partial_{i}$ within the cavity graph $\mathcal{G}_{N-1}^{(i)}$. We have omitted the normalization constant in Eq. (19), as this quantity can be fixed at the end of the calculation.
At this point, we invoke the main assumption of the cavity method on sparse random graphs Mézard and Parisi (2001); Metz et al. (2010). Since networks generated from the configuration model become locally tree-like for $N\gg 1$, the function $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}})$ factorizes as follows
$\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}})=\prod_{j\in\partial_{i}}\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}),$ (20)
leading to the following expression for $\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})$,
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}$
$\times\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j})$
$\times\prod_{j\in\partial_{i}^{\rm out}}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{ji}\rho_{i}}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}).$ (21)
In the terminology of the message-passing algorithm Mézard and Montanari (2009), $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j})$ denotes the message that propagates from node $j$ to $i$ along the directed edge $j\rightarrow i$. According to Eq. (21), the local marginal $\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})$ is determined by messages arriving at node $i$ through both ingoing and outgoing edges. However, the infection probability $\rho_{i}$ is solely determined by the in-neighbourhood $\partial_{i}^{\rm in}$, so it is sensible to assume that the messages from $\partial_{i}^{\rm out}$ contribute with a constant in Eq. (21). While we could not demonstrate this fact in full generality, making this assumption leads to results that are fully consistent with the exact solution for the dynamics of the SIS model Metz (2025).
Thus, by integrating Eq. (21) over $\hat{\rho}_{i}$ and using Eq. (17), we find
$\mathcal{P}_{N,i}(\rho)$ $=\frac{1}{Z_{N,i}}\int\limits_{0}^{1}\Bigg{(}\prod_{j\in\partial_{i}^{\rm in}}d\rho_{j}\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\Bigg{)}$
$\times\delta\Bigg{[}(1-\rho)\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}-\rho\Bigg{]},$ (22)
where $Z_{N,i}$ ensures that $\mathcal{P}_{N,i}(\rho)$ is normalized. Hence, for large $N$, $\mathcal{P}_{N,i}(\rho)$ follows from the local marginals $\{\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\}_{j\in\partial_{i}^{\rm in}}$ on $\mathcal{G}_{N-1}^{(i)}$. To determine the cavity marginals $\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})$, we follow exactly the same reasoning as explained above, which leads to the cavity equations
$\mathcal{P}_{N-1,i}^{(l)}(\rho)$ $=\frac{1}{Z_{N-1,i}^{(l)}}\int\limits_{0}^{1}\Bigg{(}\prod_{j\in\partial_{i}^{\rm in}}d\rho_{j}\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\Bigg{)}$
$\times\delta\Bigg{[}(1-\rho)\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}-\rho\Bigg{]},$ (23)
with $l\in\partial_{i}^{\rm out}$. The constant $Z_{N-1,i}^{(l)}$ normalizes $\mathcal{P}_{N-1,i}^{(l)}(\rho)$.
The solutions of Eqs. (22) and (23) provide accurate approximations for the local marginals $\{\mathcal{P}_{N,i}(\rho)\}_{i=1,\dots,N}$ on single network instances with large $N$ and a locally tree-like structure. Equation (23) becomes asymptotically exact as $N\rightarrow\infty$. In this limit, we introduce the ensemble-averaged quantities
$\mathcal{P}(\rho)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\mathcal{P}_{N,j}(\rho)$ (24)
and
$\mathcal{P}_{\rm edg}(\rho)=\lim_{N\rightarrow\infty}\frac{1}{Nc}\sum_{j=1}^{N}\sum_{l\in\partial_{j}^{\rm out}}\mathcal{P}_{N-1,j}^{(l)}(\rho).$ (25)
Assuming that in Eq. (22) the numerator and denominator converge independently to their ensemble-averaged values as $N\rightarrow\infty$, we conclude that $\mathcal{P}(\rho)$ is determined by
$\mathcal{P}(\rho)$ $=\frac{1}{Z}\sum_{k,\ell=0}^{\infty}p_{k\ell}\Bigg{(}\prod_{j=1}^{k}\int\limits_{0}^{1}d\rho_{j}\mathcal{P}_{\rm edg}(\rho_{j})\int\limits_{0}^{\infty}dx_{j}P_{\lambda}(x_{j})\Bigg{)}$
$\times\delta\Bigg{[}(1-\rho)\sum_{j=1}^{k}x_{j}\rho_{j}-\rho\Bigg{]}.$ (26)
Analogously, $\mathcal{P}_{\rm edg}(\rho)$ fulfills the self-consistent equation
$\mathcal{P}_{\rm edg}(\rho)$ $=\frac{1}{Z_{\rm edg}}\sum_{k,\ell=0}^{\infty}\frac{\ell p_{k\ell}}{c}\Bigg{(}\prod_{j=1}^{k}\int\limits_{0}^{1}d\rho_{j}\mathcal{P}_{\rm edg}(\rho_{j})\Bigg{)}$
$\times\Bigg{(}\prod_{j=1}^{k}\int\limits_{0}^{\infty}dx_{j}P_{\lambda}(x_{j})\Bigg{)}\delta\Bigg{[}(1-\rho)\sum_{j=1}^{k}x_{j}\rho_{j}-\rho\Bigg{]}.$ (27)
The constants $Z$ and $Z_{\rm edg}$ denote the corresponding normalization factors. Equations (26) and (27) are valid for networks with an arbitrary joint distribution $p_{k\ell}$ of indegrees and outdegrees. Setting $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$, we recover Eq. (11).
IV Results
In this section, we determine the phase diagram of the SIS model in the limit $N\rightarrow\infty$ by combining analytic results from random matrix theory Neri and Metz (2016); Lucas Metz et al. (2019); Metz and Neri (2021) with numerical solutions of Eq. (11).
IV.1 Linear stability analysis
First, we perform a linear stability analysis of the disease-free fixed-point $\rho_{i}=0$ ($i=1,\dots,N$) that characterizes the absorbing phase. The linearized form of Eq. (2) is given by
$\frac{d\bm{\rho}}{dt}=\left(\bm{A}-\bm{I}\right)\bm{\rho}(t),$ (28)
where $\bm{\rho}(t)=(\rho_{1}(t),\dots,\rho_{N}(t))^{T}$, and $\bm{I}$ is the $N\times N$ identity matrix. The solution of the above equation determines whether perturbations of the trivial fixed-point decay to zero or grow in time. By introducing the right $\{\bm{R}_{\alpha}\}_{\alpha=1,\dots,N}$ and left $\{\bm{L}_{\alpha}\}_{\alpha=1,\dots,N}$ eigenvectors of the asymmetric matrix $\bm{A}$,
$\bm{A}\bm{R}_{\alpha}=\Lambda_{\alpha}\bm{R}_{\alpha}\quad\bm{L}_{\alpha}\bm{A}=\Lambda_{\alpha}\bm{L}_{\alpha},$ (29)
the solution of Eq. (28) reads
$\bm{\rho}(t)=\sum_{\alpha=1}^{N}\left[\bm{L}_{\alpha}\,\bm{\rho}(0)\right]e^{(\Lambda_{\alpha}-1)t}\bm{R}_{\alpha},$ (30)
where $\Lambda_{1},\dots,\Lambda_{N}$ are the eigenvalues of $\bm{A}$. By ordering the eigenvalues according to their real parts, ${\rm Re}\Lambda_{1}\geq{\rm Re}\Lambda_{2}\geq\dots\geq{\rm Re}\Lambda_{N}$, we conclude that $|\bm{\rho}(t)|$ decays to zero for $t\rightarrow\infty$ provided
${\rm Re}\Lambda_{1}<1.$ (31)
The above condition determines the linear stability of the disease-free fixed-point.
In a series of previous works Neri and Metz (2016); Lucas Metz et al. (2019); Metz and Neri (2021), the leading eigenvalue $\Lambda_{1}$ and the statistical properties of the corresponding right eigenvector $\bm{R}_{1}$ of sparse directed networks have been computed in the limit $N\rightarrow\infty$. For $c>1$, the spectral density of $\bm{A}$ has a continuous component Neri and Metz (2020) and the spectral gap $|\Lambda_{1}-\Lambda_{2}|$ is finite if $c>c_{\rm gap}$, where
$c_{\rm gap}=1+\sigma^{2}/\lambda^{2}.$ (32)
For $c>c_{\rm gap}$, $\Lambda_{1}=c\lambda$ is an outlier Neri and Metz (2016), and the trivial solution is stable if $c<\lambda^{-1}$. For $c\leq c_{\rm gap}$, the spectral gap is zero, $\Lambda_{1}\in\mathbb{R}$ belongs to the boundary of the continuous spectrum, and the trivial fixed-point is stable for
$c<\left(\sigma^{2}+\lambda^{2}\right)^{-1}.$ (33)
These results for the stability analysis are universal, as they are independent of higher moments of the distributions $p_{k\ell}$ and $P_{\lambda}(x)$ characterizing the network structure.
As soon as the trivial fixed-point becomes unstable, the leading eigenvector $\bm{R}_{1}$ plays a crucial role for the dynamics. The moments of the real-valued eigenvector components $\{R_{1,i}\}_{i=1}^{N}$ are defined as
$\langle R_{1}^{n}\rangle=\int\limits_{-\infty}^{\infty}dr\,p_{R}(r)r^{n},$ (34)
where the distribution $p_{R}(r)$ reads
$p_{R}(r)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum\limits_{i=1}^{N}\delta(r-R_{1,i}).$ (35)
The moments of $p_{R}(r)$ characterize the fluctuations of the eigenvector components, allowing to study localization phenomena. In reference Metz and Neri (2021), the first moments of $p_{R}(r)$ have been analytically computed for sparse directed networks, unveiling a localization transition as a function of the network parameters.
IV.2 Phase diagram
Before discussing how the epidemic threshold depends on the network parameters, we compare the solutions of Eq. (11), valid for $N\rightarrow\infty$, with results obtained from the fixed-point Eq. (8) on finite-sized networks. Figure 1 shows numerical results for the prevalence $\langle\rho\rangle$ and the full distribution $\mathcal{P}(\rho)$. In both cases, the agreement between the solutions of Eq. (11) and those obtained from Eq. (8) is excellent. The distribution $\mathcal{P}(\rho)$ features a Dirac-$\delta$ at $\rho=0$, reflecting a finite fraction of nodes with zero indegree. While both approaches rely on numerical computations, an important advantage of Eq. (11) over Eq. (8) is that the former does not require the use of sophisticated algorithms to sample networks from the configuration model, since it depends on the network structure only through $p_{{\rm in},k}$ and $P_{\lambda}(x)$.
Figure 1: Comparison between the solutions of Eq. (11) (solid lines) with the fixed-point solutions of Eq. (8) (symbols) for directed networks with a $\Gamma$-distribution of infection rates and a Poisson degree distribution. The fixed-point solutions are derived from an ensemble of 5050 networks with $N=10^{5}$ nodes, while the population dynamics results are obtained from 55 independent runs of the algorithm with $M=10^{6}$ stochastic variables. (a) The prevalence $\langle\rho\rangle$ as a function of the mean infection rate $\lambda$ for standard deviation $\sigma=0.2$ of the infection rates. (b) and (c): distribution $\mathcal{P}(\rho)$ of the infection probabilities for $\lambda=1/2$, $\sigma=0.2$, and two different $c$.
In figure 2, we present the phase diagram of the SIS model on directed networks in terms of $(\sigma,c)$. The model exhibits an absorbing phase, where $\mathcal{P}(\rho)=\delta(\rho)$, and an endemic phase, characterized by a stationary distribution $\mathcal{P}(\rho)$ with nonzero prevalence $\langle\rho\rangle$. The average degree $c=\lambda^{-1}$ above which the epidemic spreads to a finite fraction of the population is determined solely by the mean infection rate and is independent of other network properties. If $(\sigma,c)$ lies above the dashed line in the phase diagram, the spectral gap $|\Lambda_{1}-\Lambda_{2}|$ is finite; otherwise, $|\Lambda_{1}-\Lambda_{2}|=0$. For $\sigma<\sigma_{*}$, the transition line that delimits the absorbing phase follows from the linear stability analysis of the disease-free fixed-point. For $\sigma>\sigma_{*}$, the transition line is obtained by numerically solving Eq. (11) and monitoring the prevalence $\langle\rho\rangle$. The colour scale in figure 2 quantifies the inverse participation ratio of the infection probabilities, which will be discussed in the next subsection.
The linear stability analysis identifies the leading eigenvector $\bm{R}_{1}$ as responsible for destabilizing the absorbing phase. By the Perron-Frobenius theorem Horn and Johnson (2012), the components of $\{R_{1,i}\}_{i=1}^{N}$ are non-negative. Above the dashed line in figure 2, $\bm{R}_{1}$ is associated with an outlier eigenvalue, characterized by $\langle R_{1}\rangle>0$ Metz and Neri (2021). This “ferromagnetic” mode triggers the onset of the endemic phase. Below the dashed line, the leading eigenvalue lies at the boundary of the continuous spectrum, where $\langle R_{1}\rangle=0$ Metz and Neri (2021). Combined with the constraint $R_{1,i}\geq 0$, this suggests that $R_{1,i}=0$ with probability one in the limit $N\rightarrow\infty$. This mode is unable to destabilize the absorbing state $\mathcal{P}(\rho)=\delta(\rho)$, which explains the absence of a transition to the endemic phase for $c<\lambda^{-1}$, in contrast with the prediction of Eq. (33). Hence, the phase transition for $\sigma>\sigma_{*}$ in figure 2 is not governed by the leading eigenpair of the contact network.
Figure 2: Phase diagram of the SIS model on directed networks in terms of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates (the mean infection rate is $\lambda=1/2$). The indegrees follow a Poisson distribution, while the infection rates follow a $\Gamma$-distribution. The model exhibits an endemic phase ($\langle\rho\rangle>0$) and an absorbing phase ($\langle\rho\rangle=0$). The standard deviations at the dots are, respectively, $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$. For $\sigma>\sigma_{*}$, the critical line is obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables. The colour scale shows the inverse participation, Eq. (38), which quantifies the spatial fluctuations of the infection probabilities. The fourth moment $\mu_{4}$ of the infection rates is defined in Eq. (42).
Figure 3: Prevalence $\langle\rho\rangle$ as a function of the mean degree $c$ for different distributions of infection rates and indegrees in the regime $\sigma<\sigma_{*}$ (see the main text). The infection rates have mean $\lambda=1/2$ and standard deviation $\sigma=0.2$. The results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{5}$ stochastic variables. The inset shows the prevalence near $c=\lambda^{-1}$ in logarithmic scale. The colours in the inset correspond to the same distributions as in the main panel.
Figure 4: Prevalence $\langle\rho\rangle$ as a function of the standard deviation $\sigma$ of the infection rates for different distributions of indegrees and infection rates in the regime $\sigma>\sigma_{*}$ (see the main text). The average indegree is $c\simeq 2.7$, and the infection rates have mean $\lambda=1/2$. For power-law distributed indegrees, the smallest indegree is $k_{\rm min}=2$. The results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables.
We now discuss the role of network heterogeneities on the prevalence $\langle\rho\rangle$ and the epidemic threshold. Figure 3 shows $\langle\rho\rangle$ as a function of $c$ for different distributions $p_{{\rm in},k}$ and $P_{\lambda}(x=\lambda_{ij})$ in the regime $\sigma<\sigma_{*}$. The prevalence vanishes as $\langle\rho\rangle\simeq c-\lambda^{-1}$ for $0<c-\lambda^{-1}\ll 1$, consistent with the mean-field critical exponent of directed percolation Ódor (2004). The epidemic threshold as well as the critical behaviour of $\langle\rho\rangle$ are both independent of the distributions $p_{{\rm in},k}$ and $P_{\lambda}(x)$, confirming the universality of the transition at $c=\lambda^{-1}$.
In the regime $\sigma>\sigma_{*}$, the stationary behaviour becomes highly sensitive to the shape of the distribution $P_{\lambda}(x)$ of infection rates. As shown in figure 4, when $\lambda_{ij}$ follows a $\Gamma$-distribution, the prevalence drops to zero at sufficiently large $\sigma$, due to the large fraction of small infection rates caused by the divergence of $P_{\lambda,{\rm g}}(x=0)$ (see Eq. (4)). In contrast, when $\lambda_{ij}$ follows a Pareto distribution, in which the smallest infection rate is $x_{0}>0$, the endemic state persists even for strong fluctuations of $\lambda_{ij}$, and $\langle\rho\rangle$ saturates at a finite value. Taken together, figures 3 and 4 show that the epidemic threshold is independent of the indegree distribution $p_{{\rm in},k}$.
IV.3 Localization of epidemic spreading
In this section we present results for higher moments of $\mathcal{P}(\rho)$, which characterize the spatial fluctuations of the infection probabilities and the localization of the epidemics. We also discuss the connection between the moments of $\mathcal{P}(\rho)$ and those of the distribution of the leading eigenvector $\bm{R}_{1}$.
The inverse participation ratio (IPR) is a standard probe of spatial localization in disordered systems Mirlin (2000); Goltsev et al. (2012); Metz and Neri (2021). Following Metz and Neri (2021), the IPR of the fixed-point vector $\bm{\rho}=(\rho_{1},\dots,\rho_{N})^{T}$ is defined as follows
$\mathcal{I}_{N}^{(4)}(\bm{\rho})=\dfrac{N\sum_{i=1}^{N}\rho_{i}^{4}}{\left(\sum_{i=1}^{N}\rho_{i}^{2}\right)^{2}},$ (36)
while the dimensionless second moment reads
$\mathcal{I}_{N}^{(2)}(\bm{\rho})=\dfrac{N\sum_{i=1}^{N}\rho_{i}^{2}}{\left(\sum_{i=1}^{N}\rho_{i}\right)^{2}}.$ (37)
If the network has a finite number of nodes with nonzero infection probabilities, $\mathcal{I}_{N}^{(4)}(\bm{\rho})$ and $\mathcal{I}_{N}^{(2)}(\bm{\rho})$ both scale linearly with $N$ and the vector $\bm{\rho}$ is localized. If the infection probabilities are nonzero on an extensive number of nodes, $\bm{\rho}$ is delocalized or extended, implying that $\mathcal{I}_{N}^{(4)}(\bm{\rho})$ and $\mathcal{I}_{N}^{(2)}(\bm{\rho})$ are of order $\mathcal{O}(N^{0})$. Here, we study the behaviour of these quantities strictly in the limit $N\rightarrow\infty$ by numerically solving Eq. (11) for $\mathcal{P}(\rho)$.
In the endemic phase, the disease infects a finite fraction of individuals and the above parameters converge to
$\mathcal{I}^{(4)}(\rho)=\lim_{N\rightarrow\infty}\mathcal{I}_{N}^{(4)}(\bm{\rho})=\dfrac{\langle\rho^{4}\rangle}{\langle\rho^{2}\rangle^{2}}$ (38)
and
$\mathcal{I}^{(2)}(\rho)=\lim_{N\rightarrow\infty}\mathcal{I}_{N}^{(2)}(\bm{\rho})=\dfrac{\langle\rho^{2}\rangle}{\langle\rho\rangle^{2}}.$ (39)
In figure 2, we quantify the fluctuations of the infection probabilities in the endemic phase by displaying $\mathcal{I}^{(4)}(\rho)$ in a colour scale. While the IPR can increase by several orders of magnitude near the phase transition, $\mathcal{I}^{(4)}(\rho)$ remains finite, indicating that the state vector $\bm{\rho}$ is delocalized in the endemic phase.
Nevertheless, the prevalence $\langle\rho\rangle$ vanishes continuously as we approach the critical line, suggesting that the disease may become localized near the epidemic threshold. To examine this scenario in more detail, we focus on the regime $\sigma<\sigma_{*}$, where the epidemic threshold is known analytically. Moreover, for $0<c-\lambda^{-1}\ll 1$ ($\sigma<\sigma_{*}$), $\bm{\rho}$ is governed by the leading eigenvector $\bm{R}_{1}$, responsible for destabilizing the absorbing phase. The moments of $\bm{R}_{1}$ have been analytically computed for directed complex networks in the limit $N\rightarrow\infty$ Metz and Neri (2021).
Figure 5: Dimensionless second moment $\mathcal{I}^{(2)}(\rho)$ [Eq. (38)] and inverse participation ratio $\mathcal{I}^{(4)}(\rho)$ [Eq. (39)] as functions of the standard deviation $\sigma$ of the infection rates for different mean degrees $c$ close to the epidemic threshold $c=\lambda^{-1}$. These results are for directed networks with Poisson indegrees and a $\Gamma$-distribution of infection rates with mean $\lambda=1/2$. Symbols represent numerical results obtained from Eq. (11) using the population dynamics algorithm with $M=10^{6}$ (vertical bars indicate the standard deviation of the mean computed over 1010 independent runs). Solid lines are analytic predictions derived from the moments of the leading eigenvector [Eqs. (40) and (41)].
For $\sigma<\sigma_{*}$, the leading eigenvalue is an outlier and the ratio $\mathcal{I}^{(2)}(R_{1})$ of the first two moments of $\bm{R}_{1}$ for a Poisson indegree distribution fulfills Metz and Pérez Castillo (2016); Lucas Metz et al. (2019); Neri and Metz (2020)
$\mathcal{I}^{(2)}(R_{1})=\dfrac{c}{c-1-\sigma^{2}/\lambda^{2}},$ (40)
while the IPR of $\bm{R}_{1}$ is given by
$\mathcal{I}^{(4)}(R_{1})=\dfrac{c^{3}\left(3\mu_{3}+4c\mu_{2}+c^{2}\right)-2c\mu_{2}^{2}\left(c^{2}+3\mu_{3}\right)}{\left(c^{3}-\mu_{4}\right)\left(c^{2}-\mu_{3}\right)},$ (41)
where we defined the dimensionless moments
$\mu_{n}=\lambda^{-n}\int_{0}^{\infty}dxP_{\lambda}(x)x^{n}$ (42)
of the infection rates. Depending on the network parameters, Eqs. (40) and (41) diverge due to vanishing denominators Metz and Neri (2021). In particular, $\mathcal{I}^{(4)}(R_{1})$ diverges at $c=\mu_{4}^{\frac{1}{3}}$. This naturally raises the question of whether $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ exhibit a similar behaviour near $c=\lambda^{-1}$.
In figure 5, we show $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ as functions of $\sigma$ near the epidemic threshold. As a comparison, this figure also displays Eqs. (40) and (41) at $c=\lambda^{-1}$, demonstrating that $\mathcal{I}^{(2)}(R_{1})$ and $\mathcal{I}^{(4)}(R_{1})$ diverge at $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$, respectively, for the $\Gamma$-distribution of infection rates. The results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ are overall consistent with the analytic predictions obtained from the moments of $\bm{R}_{1}$. The discrepancies observed in figure 5 stem from the slow convergence of the population dynamics algorithm near the thresholds $\sigma_{*}$ and $\sigma_{4}$, which makes it difficult to accurately determine the distribution $\mathcal{P}(\rho)$ from the solutions of Eq. (11). In addition, we note that $\mathcal{I}^{(4)}(\rho)$ remains bounded for $\sigma_{4}<\sigma<\sigma_{*}$, a consequence of the finite number of stochastic variables used to discretize $\mathcal{P}(\rho)$ in the population dynamics method.
Figure 6: Probability density $\mathcal{P}_{Y}(y)$ of $Y_{i}=-\ln\rho_{i}$ ($i=1,\dots,N$) near the epidemic threshold ($c-\lambda^{-1}=10^{-4}$). The directed network is characterized by a Poisson indegree distribution and a $\Gamma$-distribution of infection rates with mean $\lambda=1/2$ and varying standard deviation $\sigma$. The results are obtained from the numerical solutions of Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables. $\mathcal{P}_{Y}(y)$ exhibits an exponential decay for large $y$.
While the moments of the infection probabilities vanish as we approach the critical line, the ratios $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ diverge for sufficiently large $\sigma$, suggesting that $\bm{\rho}$ may become localized slightly above the epidemic threshold. This divergence arises from a large fraction of nodes with $\rho_{i}\simeq 0$ and a small fraction of nodes with comparatively larger $\rho_{i}$. To gain further insight into the fluctuations of $\rho_{i}$, we show in figure 6 the probability density $\mathcal{P}_{Y}(y)$ of $Y_{i}=-\ln\rho_{i}\in[0,\infty)$ near the epidemic threshold. The exponential decay $\mathcal{P}_{Y}(y)\propto e^{-Ay}$ for large $y$ implies that $\mathcal{P}(\rho)\propto\rho^{A-1}$ for $0<\rho\ll 1$, where the exponent $A$ depends on the network parameters. The power-law behaviour of $\mathcal{P}(\rho)$ near $\rho=0$ reflects strong fluctuations of $\rho_{i}$ spanning several orders of magnitude, yet confined to very small values of $\rho_{i}$.
V Final remarks
In this work, we have determined the phase diagram of the SIS model on directed complex networks within the quenched mean-field approximation. By combining random-matrix results with an analytic approach for the distribution of stationary infection probabilities, we have computed the epidemic threshold as a function of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates defining the contact network. Our results show that the SIS model exhibits a transition between the absorbing and endemic phases provided $c\geq\lambda^{-1}$, where $\lambda$ is the average infection rate.
Remarkably, the critical line is independent of the degree distribution but it depends strongly on the distribution of infection rates. When $\sigma<\sigma_{*}=\sqrt{\lambda(1-\lambda)}$, both the epidemic threshold and the critical behaviour are governed by the leading eigenpair of the contact network Metz and Neri (2021) and are independent of the infection-rate distribution. In contrast, for $\sigma>\sigma_{*}$, the fluctuations of infection rates have a pronounced effect on the critical line. While the SIS model undergoes an absorbing phase transition for a $\Gamma$-distribution of infection rates as $\sigma$ increases, it remains in the endemic phase for a Pareto distribution. This striking difference is explained by the large fraction of near-zero infection rates generated by the $\Gamma$-distribution, which suppresses the endemic state for large $\sigma$. Together, these results provide a comprehensive picture of how network heterogeneity shapes the phase diagram of the SIS model on directed networks.
We have also examined the emergence of disease localization right above the epidemic threshold Goltsev et al. (2012). Focusing on the regime where the phase transition is governed by the leading eigenpair of the contact network, we have shown that the inverse participation ratio (IPR) of the fixed-point infection probabilities diverges near the threshold for sufficiently large $\sigma$, suggesting that the disease becomes localized on a vanishing fraction of nodes. These findings are consistent with analytic predictions based on the IPR of the leading eigenvector Metz and Neri (2021). We have also computed the full probability density of infection probabilities near the critical line and found that it exhibits a large fraction of near-zero values, with strong fluctuations spanning several orders of magnitude.
We remark that our results are strictly valid in the limit $N\rightarrow\infty$, as they are primarily based on the numerical solutions of Eq. (11). Consequently, our approach does not address how the infection probabilities or the number of potentially infected nodes scale with the system size, which could in principle clarify, for instance, why the dimensionless second moment remains finite as the IPR diverges. A detailed analysis of this issue, and its relation to the finite-size scaling of the infection probabilities, is an interesting direction for future work.
From a methodological perspective, we have introduced a general analytic framework to obtain an equation for the distribution of fixed-point states in the limit $N\rightarrow\infty$. The method is not restricted to the SIS model and can be extended to study the non-equilibrium fixed-points of other dynamical systems on directed networks, including the Kuramoto model Acebrón et al. (2005) and firing-rate models of neural networks Sompolinsky et al. (1988). Moreover, since the approach is based on the cavity method, it can also incorporate other important network features, such as short loops and degree-degree correlations, opening promising avenues of future research.
Acknowledgements.
V. B. M. acknowledges fellowships from CNPq/Brazil and BIC-UFRGS. F. L. M. acknowledges support from CNPq (Grant No 402487/2023-0), FAPERJ (Grant No 204.646/2024), and from ICTP through the Associates Program (2023-2028)