### v2

# Epidemic threshold and localization of the SIS model on directed complex networks

Abstract
We study the susceptible-infected-susceptible (SIS) model on directed complex networks within the quenched mean-field approximation. Combining results from random matrix theory with an analytic approach to the distribution of fixed-point infection probabilities, we derive the phase diagram and show that the model exhibits a nonequilibrium phase transition between the absorbing and endemic phases for $c\geq\lambda^{-1}$, where $c$ is the mean degree and $\lambda$ the average infection rate. Interestingly, the critical line is independent of the degree distribution but is highly sensitive to the form of the infection-rate distribution. We further show that the inverse participation ratio of infection probabilities diverges near the epidemic threshold, indicating that the disease may become localized on a small fraction of nodes. These results provide a systematic characterization of how network heterogeneities shape epidemic spreading on directed contact networks within the quenched mean-field approximation.
preprint: APS/123-QED
I Introduction
The susceptible-infected-susceptible (SIS) model on networks provides a minimal yet powerful framework to investigate the interplay between epidemic spreading and the structure of the underlying contact network Kiss et al. (2017). In the SIS model, each node (individual) can be in either a susceptible or infected state. A susceptible node $i$ can become infected by a neighbour $j$ at a rate $\lambda_{ij}$, typically assumed to be uniform, $\lambda_{ij}=\lambda$. An infected node, in turn, becomes healthy at a rate conventionally set to unity. The average fraction of infected individuals, referred to as the prevalence, is the natural order-parameter for characterizing the spread of the epidemic. In the limit of an infinitely large number $N$ of individuals, the SIS model may exhibit a nonequilibrium phase transition between an absorbing phase Pastor-Satorras et al. (2015), in which the epidemic dies out, and an endemic phase, characterized by a stationary state with a nonzero prevalence.
A central problem in the study of epidemic spreading is to understand how the network structure influences the epidemic threshold $\lambda_{\rm c}$ Chakrabarti et al. (2008); Chatterjee and Durrett (2009); Castellano and Pastor-Satorras (2010); Ferreira et al. (2012); Goltsev et al. (2012); Li et al. (2013); Van Mieghem and van de Bovenkamp (2013); Kwon and Kim (2013), which separates the absorbing phase ($\lambda\leq\lambda_{\rm c}$) from the endemic phase ($\lambda>\lambda_{\rm c}$). The quenched mean-field (QMF) theory Ferreira et al. (2012); Mata and Ferreira (2013); Pastor-Satorras et al. (2015) provides an effective approximation for the SIS model by neglecting dynamical correlations between neighboring nodes Mata and Ferreira (2013), leading to a set of coupled dynamical equations for the single-node infection probabilities. Within this framework, the epidemic threshold for a network of size $N$ is given by Ferreira et al. (2012); Van Mieghem and van de Bovenkamp (2013)
$\lambda_{\rm c}(N)=1/\Lambda_{1}(N),$ (1)
where $\Lambda_{1}(N)>0$ denotes the leading eigenvalue (spectral radius) of the network adjacency matrix Newman (2010). Beyond its role in QMF theory, Eq. (1) provides a rigorous lower bound for the exact epidemic threshold on any given network Van Mieghem and van de Bovenkamp (2013). The QMF approximation can be improved by incorporating pairwise dynamical correlations, leading to the so-called pair quenched mean-field theory Mata and Ferreira (2013), which yields more accurate predictions for the epidemic threshold compared to the individual-based QMF theory.
The spectral properties of the adjacency matrix thus play a fundamental role in the dynamics of the SIS model. For undirected networks, the epidemic threshold follows from well-established analytic results for the leading eigenvalue $\Lambda_{1}(N)$ Krivelevich and Sudakov (2003); Chung et al. (2003). In such networks, the expected value of $\Lambda_{1}(N)$ typically scales with the maximum degree. Consequently, for degree distributions with unbounded support, the epidemic threshold vanishes in the limit $N\rightarrow\infty$ Castellano and Pastor-Satorras (2010); Ferreira et al. (2012). This prediction of the QMF theory is fully consistent with rigorous results for the SIS model Chatterjee and Durrett (2009), which prove the absence of an absorbing phase for any nonzero infection rate in undirected networks.
Another spectral property relevant to the dynamics of the SIS model is the spatial localization of the leading eigenvector associated with $\Lambda_{1}(N)$ Goltsev et al. (2012); Pastor-Satorras and Castellano (2018). A strongly localized eigenvector has a finite number of nonzero components as $N\rightarrow\infty$ Abou-Chacra et al. (1973); Mirlin (2000), whereas an extended eigenvector is characterized by an extensive number of nonzero components. Localization effects are expected to be particularly relevant near the epidemic threshold, where the leading eigenvector closely approximates the stationary endemic state, providing insights into network-based immunization strategies that target influential nodes Pastor-Satorras and Castellano (2018). In undirected networks, the localization of the leading eigenvector has been studied through the inverse participation ratio (IPR) Goltsev et al. (2012); Pastor-Satorras and Castellano (2016, 2018); Liu and Van Mieghem (2019), defined in terms of the fourth moment of the eigenvector components. Numerical results for the IPR indicate that, in finite-size undirected networks, the leading eigenvector is localized on a vanishing fraction of nodes Goltsev et al. (2012); Pastor-Satorras and Castellano (2016, 2018). In particular, for degree distributions that decay sufficiently fast, this localization occurs at the hub with the largest degree Goltsev et al. (2012); Pastor-Satorras and Castellano (2016).
In contrast to undirected networks, the SIS model on directed contact networks, where the coupling strengths or infection rates $\lambda_{ij}$ are unidirectional  [^1] In a directed contact network, if node $j$ infects node $i$ with rate $\lambda_{ij}$, the opposite rate $\lambda_{ji}$ is zero., remains poorly understood. Spreading models on directed networks provide a natural framework for studying the transmission of computer viruses through email networks Newman et al. (2002); Ebel et al. (2002), the diffusion of information in social networks Krueger et al. (2024), and the transmission of diseases between patients and health care workers Meyers et al. (2006). Only a few works have examined how directionality impacts the SIS model Li et al. (2013); Kwon and Kim (2013). Numerical studies on finite networks with both directed and bidirected edges have shown that $\lambda_{\rm c}(N)$ increases with the fraction of directed links Li et al. (2013). In addition, the SIS model has been studied on directed networks using heterogeneous mean-field theory Kwon and Kim (2013), which does not fully capture the underlying network structure. As a result, even basic problems, such as how the directed network structure affects the epidemic threshold, remain unresolved.
In this work, we study the QMF theory of the SIS model on directed networks with arbitrary distributions of degrees and infection rates. We determine the epidemic threshold $\lambda_{\rm c}$ in the limit $N\rightarrow\infty$, showing that this model undergoes an absorbing phase transition as a function of the network structure. Our calculation of $\lambda_{\rm c}$ partially relies on random-matrix analytic results for the leading eigenpair of directed weighted networks Metz and Pérez Castillo (2016); Lucas Metz et al. (2019); Metz and Neri (2021). In contrast with undirected networks, the leading eigenvalue of directed networks remains finite as $N\rightarrow\infty$ Metz and Pérez Castillo (2016), even when the degree distribution has unbounded support.
We show that the epidemic threshold is determined by the leading eigenvalue only when the network parameters are such that the gap between the leading and the subleading eigenvalue remains finite as $N\rightarrow\infty$. With the main goal of computing $\lambda_{\rm c}$ for arbitrary combinations of network parameters, we derive an exact equation for the full distribution of the stationary infection probabilities in the limit $N\rightarrow\infty$, using the cavity method from spin-glass theory Mézard and Parisi (2001); Metz et al. (2010). The numerical solutions of this equation yield both the epidemic threshold and the IPR as functions of the network structure. We find that $\lambda_{\rm c}$ is independent of the degree distribution, whereas the shape of the infection-rate distribution strongly influences the epidemic threshold, in particular for large fluctuations in the infection rates. Furthermore, we show that the IPR diverges near the epidemic threshold as $N\rightarrow\infty$, in agreement with analytic predictions derived from the moments of the leading eigenvector Metz and Neri (2021). This divergence arises from a large fraction of nodes with infection probabilities that strongly fluctuate near zero, indicating that the disease becomes localized on a vanishing fraction of network nodes.
The paper is organized as follows. In the next section, we introduce the SIS model on directed complex networks within the QMF approximation. In section III, we derive the equation for the stationary distribution of infection probabilities using the cavity method. The results for the epidemic threshold and the IPR as functions of the network parameters are discussed in section IV. Finally, we present a summary of our findings and concluding remarks in section V.
II The SIS model on directed networks
We consider the SIS model in the quenched mean-field approximation Pastor-Satorras et al. (2015); Mata and Ferreira (2013). The probability $\rho_{i}(t)$ ($i=1,\dots,N$) that a node $i$ is infected at time $t$ evolves according to
$\frac{d\rho_{i}}{dt}=-\rho_{i}(t)+\left[1-\rho_{i}(t)\right]\sum_{j=1}^{N}A_{ij}\rho_{j}(t),$ (2)
where the elements $\{A_{ij}\}_{i,j=1,\dots,N}$ of the $N\times N$ weighted adjacency matrix $\bm{A}$ have the form $A_{ij}=C_{ij}\lambda_{ij}$, with $C_{ij}\in\{0,1\}$ and $\lambda_{ij}>0$. The binary variables $\{C_{ij}\}_{i,j=1,\dots,N}$ specify the structure of the contact network: if $C_{ij}=1$, there is a directed link $j\rightarrow i$ from node $j$ to $i$, whereas $C_{ij}=0$ otherwise. The diagonal elements $\{C_{ii}\}_{i=1,\dots,N}$ are zero. The indegree $K_{i}$ (outdegree $L_{i}$) of node $i$ is given by $K_{i}=\sum_{j=1}^{N}C_{ij}$ ($L_{i}=\sum_{j=1}^{N}C_{ji}$). We consider directed random networks generated from the configuration model Newman et al. (2001); Fosdick et al. (2018), where the degree sequences $K_{1},\dots,K_{N}$ and $L_{1},\dots,L_{N}$ are independent and identically distributed random variables drawn from $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$, with $p_{{\rm in},k}$ and $p_{{\rm out},\ell}$ denoting, respectively, the indegree and outdegree distributions. The degree sequences fulfill the constraint $\sum_{i=1}^{N}K_{i}=\sum_{i=1}^{N}L_{i}$, and the average degree $c$ reads
$c=\sum_{k=0}^{\infty}k\,p_{{\rm in},k}=\sum_{\ell=0}^{\infty}\ell\,p_{{\rm out},\ell}.$ (3)
We will present results for three examples of degree distributions: Poisson, geometric and power-law Newman et al. (2001).
The coefficient $\lambda_{ij}>0$ denotes the directed coupling strength $j\rightarrow i$ or the pairwise infection rate at which node $j$ infects $i$. We assume that $\{\lambda_{ij}\}_{i,j=1,\dots,N}$ are independent and identically distributed random variables sampled from a distribution $P_{\lambda}(x=\lambda_{ij})$ with mean $\lambda$ and variance $\sigma^{2}$. We will present results for two choices of $P_{\lambda}(x)$. The first one is the $\Gamma$-distribution
$P_{\lambda,{\rm g}}(x)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-\beta x}\Theta(x),$ (4)
with parameters
$\alpha=\lambda^{2}/\sigma^{2}\quad\text{and}\quad\beta=\lambda/\sigma^{2},$ (5)
and with $\Theta(x)$ representing the Heaviside step function. The second example is the Pareto distribution with finite variance,
$P_{\lambda,{\rm p}}(x)=\frac{\gamma x_{0}^{\gamma}}{x^{\gamma+1}}\Theta\left(x-x_{0}\right),$ (6)
where
$\gamma=1+\sqrt{1+\frac{\lambda^{2}}{\sigma^{2}}}>2\quad\text{and}\quad x_{0}=\left(\frac{\sqrt{1+\frac{\lambda^{2}}{\sigma^{2}}}}{1+\sqrt{1+\frac{\lambda^{2}}{\sigma^{2}}}}\right)\lambda.$ (7)
Note that both distributions are solely parametrized in terms of $\lambda$ and $\sigma$. For $\sigma=0$, the infection rates are uniform ($\lambda_{ij}=\lambda$).
The network ensemble is fully specified by the distributions $p_{k\ell}$ and $P_{\lambda}(x)$. For $c>1$, the network contains a giant strongly connected component Kryven (2016), ensuring that the spectrum of $\bm{A}$ has a continuous component in the limit $N\rightarrow\infty$ Neri and Metz (2020). This is the interesting regime where the nodes strongly interact with each other and the disease can eventually infect a finite fraction of individuals.
Our goal is to understand how the structure of the contact network influences the stationary states of the model. Setting $\frac{d\rho_{i}}{dt}=0$ in Eq. (2), we obtain the fixed-point equations
$\rho_{i}=\frac{\sum_{j=1}^{N}A_{ij}\rho_{j}}{1+\sum_{j=1}^{N}A_{ij}\rho_{j}}\quad(i=1,\dots,N).$ (8)
To characterize the stationary behaviour and the phase diagram in the limit $N\rightarrow\infty$, we study the distribution
$\mathcal{P}(\rho)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\delta(\rho-\rho_{i})$ (9)
of the fixed-point infection probabilities $\rho_{1},\dots,\rho_{N}$. The moments of $\mathcal{P}(\rho)$ read
$\langle\rho^{n}\rangle=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\rho_{i}^{n}=\int_{0}^{1}d\rho\mathcal{P}(\rho)\rho^{n}.$ (10)
In particular, the prevalence $\langle\rho\rangle$ quantifies the average fraction of infected individuals.
In the next section, we develop an analytic approach, based on the cavity method of spin-glass theory Mézard and Parisi (2001); Metz et al. (2010), that yields the following self-consistent equation for the distribution $\mathcal{P}(\rho)$,
$\mathcal{P}(\rho)=\frac{1}{Z}\sum_{k=0}^{\infty}p_{{\rm in},k}\left(\prod_{j=1}^{k}\int_{0}^{1}d\rho_{j}\mathcal{P}(\rho_{j})\int_{0}^{\infty}dx_{j}P_{\lambda}(x_{j})\right)$
$\times\delta\Bigg[(1-\rho)\sum_{j=1}^{k}x_{j}\rho_{j}-\rho\Bigg],$ (11)
where the constant $Z$ ensures that $\int_{0}^{1}d\rho\mathcal{P}(\rho)=1$. This equation can be also inferred from the effective dynamics obtained in Metz (2025) using dynamical mean-field theory.
Equation (11) can be numerically solved through a Monte-Carlo iterative method known as population dynamics Mézard and Parisi (2001); Mézard and Montanari (2009); Metz and Pérez Castillo (2016); Lucas Metz et al. (2019). The method is based on the discretization of $\mathcal{P}(\rho)$ in terms of a population with $M$ stochastic variables $\rho_{1},\dots,\rho_{M}$. In the standard version of the algorithm Mézard and Parisi (2001); Mézard and Montanari (2009); Metz and Pérez Castillo (2016), each iteration proceeds by randomly selecting a member of the population and updating it according to the constraint imposed by the Dirac-$\delta$ in Eq. (11). After a sufficient number of iterations, the population $\rho_{1},\dots,\rho_{M}$ converges to a stationary profile, yielding an approximate solution of Eq. (11). In the present context, the core of the algorithm remains unchanged, but we have to include the weight arising from the normalization factor $Z$, which determines the number of elements updated in parallel at each iteration. The details of this modified version of the algorithm are discussed in Lucas Metz et al. (2019).
III The cavity approach for the distribution of infection probabilities
In this section, we present a first-principles derivation of Eq. (11) using the cavity method from spin-glass theory. The dynamic cavity method has been developed in the context of epidemic spreading to estimate the probability $\rho_{i}(t)$ that node $i$ is infected at time $t$ Lokhov et al. (2014, 2015); Shrestha et al. (2015); Ortega et al. (2022), leading to important applications such as the inference of the origin of an epidemic outbreak Lokhov et al. (2014); Altarelli et al. (2014a) and the design of optimal immunization strategies Altarelli et al. (2014b). Here, instead of following a dynamical approach, we start from the fixed-point equations for the infection probabilities $\rho_{1},\dots,\rho_{N}$ in the QMF approximation and compute their empirical distribution using the cavity method. We note that this section can be skipped by readers primarily interested in the results obtained by solving Eq. (11).
The variables $\bm{\rho}=(\rho_{1},\dots,\rho_{N})^{T}$ solve the equations
$F_{N,i}(\bm{\rho}|\bm{h})=-\rho_{i}+(1-\rho_{i})\sum_{j=1}^{N}A_{ij}\rho_{j}+h_{i}=0,$ (12)
where $i=1,\dots,N$. We have added auxiliary fields $\bm{h}=(h_{1},\dots,h_{N})^{T}$ to the fixed-point equations, which are useful to identify the physical meaning of certain quantities. These fields will be set to zero at the end of the calculation. The joint probability density of $\bm{\rho}$ can be formally written as Kurchan (1991)
$\mathcal{P}_{N}(\bm{\rho}|\bm{h})=\dfrac{\prod_{i=1}^{N}\delta\left[F_{N,i}(\bm{\rho}|\bm{h})\right]}{\int\limits_{0}^{1}\left(\prod_{i=1}^{N}d\rho_{i}\right)\prod_{i=1}^{N}\delta\left[F_{N,i}(\bm{\rho}|\bm{h})\right]}.$ (13)
Assuming that Eqs. (12) admit a unique stable solution, Eq. (13) concentrates as $N\rightarrow\infty$ on the joint distribution of infection probabilities characterizing this stable state. Starting from Eq. (13), our purpose is to compute the local marginals $\{\mathcal{P}_{N,i}(\rho_{i})\}_{i=1}^{N}$ on the network nodes by using the cavity method Metz et al. (2010). Using the Fourier transform of the Dirac-$\delta$, we write $\mathcal{P}_{N}(\bm{\rho}|\bm{h})$ as
$\mathcal{P}_{N}(\bm{\rho}|\bm{h})=\int_{-\infty}^{\infty}\left(\prod_{i=1}^{N}d\hat{\rho}_{i}\right)\gamma_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h}),$ (14)
where
$\gamma_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=\left[Z_{N}(\bm{h})\right]^{-1}\exp{\left[H_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})\right]},$ (15)
with
$H_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=i\sum_{j=1}^{N}h_{j}\hat{\rho}_{j}-i\sum_{j=1}^{N}\hat{\rho}_{j}\rho_{j}+i\sum_{l,j=1}^{N}A_{lj}\hat{\rho}_{l}(1-\rho_{l})\rho_{j}.$ (16)
[…]
From Eq. (14), the marginal $\mathcal{P}_{N,i}(\rho_{i}|h_{i})$ is given by
$\mathcal{P}_{N,i}(\rho_{i}|h_{i})=\int\limits_{-\infty}^{\infty}d\hat{\rho}_{i}\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}|h_{i}).$ (17)
Let us extract the $i$-th term from the sums in Eq. (16), namely
$H_{N}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=ih_{i}\hat{\rho}_{i}-i\hat{\rho}_{i}\rho_{i}+i\rho_{i}\sum_{j\in\partial_{i}^{\rm out}}\lambda_{ji}\hat{\rho}_{j}(1-\rho_{j})$
$+i\hat{\rho}_{i}(1-\rho_{i})\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}+H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h}).$ (18)
The function $H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})$ is defined on the cavity graph $\mathcal{G}_{N-1}^{(i)}$, which is obtained from the original graph $\mathcal{G}_{N}$ by removing node $i$, its adjacent edges, and the local field $h_{i}$. The symbol $\partial_{i}^{\rm out}$ denotes the set of nodes that receive a directed link from $i$ (the out-neighborhood of $i$), while $\partial_{i}^{\rm in}$ represents the set of nodes that have a directed link pointing to $i$ (the in-neighborhood of $i$). The full neighborhood of $i$ is defined as $\partial_{i}=\partial_{i}^{\rm out}\cup\partial_{i}^{\rm in}$. By substituting Eq. (18) in Eq. (15) and integrating the resulting expression over all variables except $(\rho_{i},\hat{\rho}_{i})$, we find
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}|h_{i})\sim e^{i\hat{\rho}_{i}h_{i}-i\hat{\rho}_{i}\rho_{i}}\int\limits_{-\infty}^{\infty}\Bigg(\prod_{j\in\partial_{i}}d\hat{\rho}_{j}\Bigg)\int\limits_{0}^{1}\Bigg(\prod_{j\in\partial_{i}}d\rho_{j}\Bigg)$
$\times\exp{\Bigg[i\rho_{i}\sum_{j\in\partial_{i}^{\rm out}}\lambda_{ji}\hat{\rho}_{j}(1-\rho_{j})+i\hat{\rho}_{i}(1-\rho_{i})\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}\Bigg]}$
$\times\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}}),$ (19)
where $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}})$ is the joint distribution defined on the neighbourhood $\partial_{i}$ within the cavity graph $\mathcal{G}_{N-1}^{(i)}$. We have omitted the normalization constant in Eq. (19), as this quantity can be fixed at the end of the calculation.
At this point, we invoke the main assumption of the cavity method on sparse random graphs Mézard and Parisi (2001); Metz et al. (2010). Since networks generated from the configuration model become locally tree-like for $N\gg 1$, the function $\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}})$ factorizes as follows
$\gamma_{N-1,\partial_{i}}^{(i)}(\rho_{\partial_{i}},\hat{\rho}_{\partial_{i}}|h_{\partial_{i}})=\prod_{j\in\partial_{i}}\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j}).$ (20)
Inserting the above expression in Eq. (19) and using Eq. (17), we find
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i}|h_{i})\sim e^{i\hat{\rho}_{i}h_{i}-i\hat{\rho}_{i}\rho_{i}}$
$\times\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\mathcal{P}_{N-1,j}^{(i)}(\rho_{j}|h_{j})$
$\times\prod_{j\in\partial_{i}^{\rm out}}\int\limits_{0}^{1}d\rho_{j}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{ji}\rho_{i}}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j}).$ (21)
In the terminology of the message-passing algorithm Mézard and Montanari (2009), $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$ denotes the message that propagates from node $j$ to $i$ along the directed edge $j\rightarrow i$. Since a node $i$ can be infected only by its in-neighbourhood $\partial_{i}^{\rm in}$, the term containing the messages $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$ originated in the out-neighbourhood $j\in\partial_{i}^{\rm out}$ of $i$ should be irrelevant and contribute with a constant in Eq. (21).
Let us now demonstrate this property. By repeating the same reasoning that led to Eq. (21), we can derive the cavity equations for $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$. For $i\in\partial_{j}^{\rm in}$, we obtain
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})\sim e^{i\hat{\rho}_{j}h_{j}-i\hat{\rho}_{j}\rho_{j}}$
$\times\prod_{l\in\partial_{j}^{\rm in}\setminus i}\,\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{jl}\rho_{l}}\,\mathcal{P}_{N-1,l}^{(j)}(\rho_{l}|h_{l})$
$\times\prod_{l\in\partial_{j}^{\rm out}}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{l}\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{l}(1-\rho_{l})\lambda_{lj}\rho_{j}}\,\gamma_{N-1,l}^{(j)}(\rho_{l},\hat{\rho}_{l}|h_{l}),$ (22)
while for $i\in\partial_{j}^{\rm out}$ we get
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})\sim e^{i\hat{\rho}_{j}h_{j}-i\hat{\rho}_{j}\rho_{j}}$
$\times\prod_{l\in\partial_{j}^{\rm in}}\,\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{j}(1-\rho_{j})\lambda_{jl}\rho_{l}}\,\mathcal{P}_{N-1,l}^{(j)}(\rho_{l}|h_{l})$
$\times\prod_{l\in\partial_{j}^{\rm out}\setminus i}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{l}\int\limits_{0}^{1}d\rho_{l}\,e^{i\hat{\rho}_{l}(1-\rho_{l})\lambda_{lj}\rho_{j}}\,\gamma_{N-1,l}^{(j)}(\rho_{l},\hat{\rho}_{l}|h_{l}),$ (23)
where $\partial_{j}^{\rm in}\setminus i$ ($\partial_{j}^{\rm out}\setminus i$) represents the set of nodes in the in-neighbourhood (out-neighbourhood) of $j$, except for node $i$. Let us focus on $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})$ with $i\in\partial_{j}^{\rm in}$, which is precisely the quantity appearing in the right hand side of Eq. (21). By setting all auxiliary fields to zero in Eq. (22), except for $h_{j}$, we obtain the interesting identity
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|h_{j})=e^{i\hat{\rho}_{j}h_{j}}\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}),$ (24)
with $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j})\equiv\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|0)$. Hence, setting all external fields to zero in Eq. (21) and using the above identity, we find
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})$
$\times\prod_{j\in\partial_{i}^{\rm out}}\int\limits_{0}^{1}d\rho_{j}\int\limits_{-\infty}^{\infty}d\hat{\rho}_{j}\,\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|(1-\rho_{j})\lambda_{ji}\rho_{i}).$ (25)
The final step of this calculation is to understand the physical meaning of $\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|(1-\rho_{j})\lambda_{ji}\rho_{i})$. This can be achieved by extracting the $j$-th term from the function $H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})$ on the cavity graph $\mathcal{G}_{N-1}^{(i)}$,
$H_{N-1}^{(i)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h})=ih_{j}\hat{\rho}_{j}-i\hat{\rho}_{j}\rho_{j}+i\rho_{j}\sum_{l\in\partial_{j}^{\rm out}}\lambda_{lj}\hat{\rho}_{l}(1-\rho_{l})$
$+i\hat{\rho}_{j}(1-\rho_{j})\sum_{l\in\partial_{j}^{\rm in}\setminus i}\lambda_{jl}\rho_{l}+H_{N-2}^{(i,j)}(\bm{\rho},\hat{\bm{\rho}}|\bm{h}),$ (26)
where $i\in\partial_{j}^{\rm in}$. If we choose $h_{j}=(1-\rho_{j})\lambda_{ji}\rho_{i}$ in the above equation, we restore the neighbourhood $\partial_{j}$ of $j$ as in the original graph. In other words, by setting the external field to $h_{j}=(1-\rho_{j})\lambda_{ji}\rho_{i}$, we add back the interaction term between $j$ and $i$, which has been deleted through the removal of $i$. Hence, for $N\gg 1$, the following relation holds
$\gamma_{N-1,j}^{(i)}(\rho_{j},\hat{\rho}_{j}|(1-\rho_{j})\lambda_{ji}\rho_{i})=\gamma_{N,j}(\rho_{j},\hat{\rho}_{j}).$ (27)
Inserting the above relation in Eq. (25) and using the normalization of $\gamma_{N}(\rho_{j},\hat{\rho}_{j})$, we obtain
$\gamma_{N,i}(\rho_{i},\hat{\rho}_{i})\sim e^{-i\hat{\rho}_{i}\rho_{i}}\prod_{j\in\partial_{i}^{\rm in}}\int\limits_{0}^{1}d\rho_{j}\,e^{i\hat{\rho}_{i}(1-\rho_{i})\lambda_{ij}\rho_{j}}\,\mathcal{P}_{N-1,j}^{(i)}(\rho_{j}),$ (28)
as expected.
Finally, by integrating Eq. (28) over $\hat{\rho}_{i}$, we find
$\mathcal{P}_{N,i}(\rho)$ $=\frac{1}{Z_{N,i}}\int\limits_{0}^{1}\Bigg(\prod_{j\in\partial_{i}^{\rm in}}d\rho_{j}\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\Bigg)$
$\times\delta\Bigg[(1-\rho)\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}-\rho\Bigg],$ (29)
where $Z_{N,i}$ ensures that $\mathcal{P}_{N,i}(\rho)$ is normalized. Hence, for large $N$, $\mathcal{P}_{N,i}(\rho)$ follows from the local marginals $\{\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\}_{j\in\partial_{i}^{\rm in}}$ on $\mathcal{G}_{N-1}^{(i)}$. To determine the cavity marginals $\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})$, we start from Eq. (23) and follow exactly the same reasoning as explained above, which leads to the cavity equations
$\mathcal{P}_{N-1,i}^{(l)}(\rho)$ $=\frac{1}{Z_{N-1,i}^{(l)}}\int\limits_{0}^{1}\Bigg(\prod_{j\in\partial_{i}^{\rm in}}d\rho_{j}\mathcal{P}_{N-1,j}^{(i)}(\rho_{j})\Bigg)$
$\times\delta\Bigg[(1-\rho)\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}-\rho\Bigg],$ (30)
with $l\in\partial_{i}^{\rm out}$. The constant $Z_{N-1,i}^{(l)}$ normalizes $\mathcal{P}_{N-1,i}^{(l)}(\rho)$. The local marginals $\mathcal{P}_{N,i}(\rho)$ and $\mathcal{P}_{N-1,i}^{(l)}(\rho)$ fulfill the same equation on a single graph instance because the quantity $\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}\rho_{j}$ is solely determined by the in-neighbourhood $\partial_{i}^{\rm in}$, while $\mathcal{P}_{N-1,i}^{(l)}(\rho)$ is obtained by removing a node $l$ from the out-neighbourhood. This is different, for instance, from the cavity approach in undirected networks, where there is no distinction between in-neighbourhood and out-neighbourhood, and the removal of a node influences the states of all its neighbours.
The solutions of Eqs. (29) and (30) provide accurate approximations for the local marginals $\{\mathcal{P}_{N,i}(\rho)\}_{i=1,\dots,N}$ on single network instances with large $N$ and a locally tree-like structure. Equation (30) becomes asymptotically exact as $N\rightarrow\infty$. In this limit, we introduce the ensemble-averaged quantities
$\mathcal{P}(\rho)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\mathcal{P}_{N,j}(\rho)$ (31)
and
$\mathcal{P}_{\rm edg}(\rho)=\lim_{N\rightarrow\infty}\frac{1}{Nc}\sum_{j=1}^{N}\sum_{l\in\partial_{j}^{\rm out}}\mathcal{P}_{N-1,j}^{(l)}(\rho).$ (32)
Assuming that in Eq. (29) the numerator and denominator converge independently to their ensemble-averaged values as $N\rightarrow\infty$, we conclude that $\mathcal{P}(\rho)$ is determined by
$\mathcal{P}(\rho)$ $=\frac{1}{Z}\sum_{k,\ell=0}^{\infty}p_{k\ell}\Bigg(\prod_{j=1}^{k}\int\limits_{0}^{1}d\rho_{j}\mathcal{P}_{\rm edg}(\rho_{j})\int\limits_{0}^{\infty}dx_{j}P_{\lambda}(x_{j})\Bigg)$
$\times\delta\Bigg[(1-\rho)\sum_{j=1}^{k}x_{j}\rho_{j}-\rho\Bigg].$ (33)
Analogously, $\mathcal{P}_{\rm edg}(\rho)$ fulfills the self-consistent equation
$\mathcal{P}_{\rm edg}(\rho)$ $=\frac{1}{Z_{\rm edg}}\sum_{k,\ell=0}^{\infty}\frac{\ell p_{k\ell}}{c}\Bigg(\prod_{j=1}^{k}\int\limits_{0}^{1}d\rho_{j}\mathcal{P}_{\rm edg}(\rho_{j})\Bigg)$
$\times\Bigg(\prod_{j=1}^{k}\int\limits_{0}^{\infty}dx_{j}P_{\lambda}(x_{j})\Bigg)\delta\Bigg[(1-\rho)\sum_{j=1}^{k}x_{j}\rho_{j}-\rho\Bigg].$ (34)
The constants $Z$ and $Z_{\rm edg}$ denote the corresponding normalization factors. Equations (33) and (34) are valid for networks with an arbitrary joint distribution $p_{k\ell}$ of indegrees and outdegrees. Substituting $p_{k\ell}=p_{{\rm in},k}p_{{\rm out},\ell}$ in Eq. (34) and summing over $l$, we conclude that $\mathcal{P}(\rho)=\mathcal{P}_{\rm edg}(\rho)$, recovering Eq. (11).
IV Results
In this section, we determine the phase diagram of the SIS model in the limit $N\rightarrow\infty$ by combining analytic results from random matrix theory Neri and Metz (2016); Lucas Metz et al. (2019); Metz and Neri (2021) with numerical solutions of Eq. (11).
IV.1 Linear stability analysis
First, we perform a linear stability analysis of the disease-free fixed-point $\rho_{i}=0$ ($i=1,\dots,N$) that characterizes the absorbing phase. The linearized form of Eq. (2) is given by
$\frac{d\bm{\rho}}{dt}=\left(\bm{A}-\bm{I}\right)\bm{\rho}(t),$ (35)
where $\bm{\rho}(t)=(\rho_{1}(t),\dots,\rho_{N}(t))^{T}$, and $\bm{I}$ is the $N\times N$ identity matrix. The solution of the above equation determines whether perturbations of the trivial fixed-point decay to zero or grow in time. By introducing the right $\{\bm{R}_{\alpha}\}_{\alpha=1,\dots,N}$ and left $\{\bm{L}_{\alpha}\}_{\alpha=1,\dots,N}$ eigenvectors of the asymmetric matrix $\bm{A}$,
$\bm{A}\bm{R}_{\alpha}=\Lambda_{\alpha}\bm{R}_{\alpha}\quad\bm{L}_{\alpha}\bm{A}=\Lambda_{\alpha}\bm{L}_{\alpha},$ (36)
the solution of Eq. (35) reads
$\bm{\rho}(t)=\sum_{\alpha=1}^{N}\left[\bm{L}_{\alpha}\,\bm{\rho}(0)\right]e^{(\Lambda_{\alpha}-1)t}\bm{R}_{\alpha},$ (37)
where $\Lambda_{1},\dots,\Lambda_{N}$ are the eigenvalues of $\bm{A}$. By ordering the eigenvalues according to their real parts, ${\rm Re}\Lambda_{1}\geq{\rm Re}\Lambda_{2}\geq\dots\geq{\rm Re}\Lambda_{N}$, we conclude that $|\bm{\rho}(t)|$ decays to zero for $t\rightarrow\infty$ provided
${\rm Re}\Lambda_{1}<1.$ (38)
The above condition determines the linear stability of the disease-free fixed-point.
In a series of previous works Neri and Metz (2016); Lucas Metz et al. (2019); Metz and Neri (2021), the leading eigenvalue $\Lambda_{1}$ and the statistical properties of the corresponding right eigenvector $\bm{R}_{1}$ of sparse directed networks have been computed in the limit $N\rightarrow\infty$. For $c>1$, the spectral density of $\bm{A}$ has a continuous component Neri and Metz (2020) and the spectral gap $|\Lambda_{1}-\Lambda_{2}|$ is finite if $c>c_{\rm gap}$, where
$c_{\rm gap}=1+\sigma^{2}/\lambda^{2}.$ (39)
For $c>c_{\rm gap}$, $\Lambda_{1}=c\lambda$ is an outlier Neri and Metz (2016), and the trivial solution is stable if $c<\lambda^{-1}$. For $c\leq c_{\rm gap}$, the spectral gap is zero, $\Lambda_{1}=\sqrt{c\left(\sigma^{2}+\lambda^{2}\right)}\in\mathbb{R}$ belongs to the boundary of the continuous spectrum, and the trivial fixed-point is stable for
$c<\left(\sigma^{2}+\lambda^{2}\right)^{-1}.$ (40)
These results for the stability analysis hold for $c>1$. The stability thresholds are universal, as they are independent of higher moments of the distributions $p_{k\ell}$ and $P_{\lambda}(x)$ characterizing the network structure.
As soon as the trivial fixed-point becomes unstable, the leading eigenvector $\bm{R}_{1}$ plays a crucial role for the dynamics. The moments of the real-valued eigenvector components $\{R_{1,i}\}_{i=1}^{N}$ are defined as
$\langle R_{1}^{n}\rangle=\int\limits_{-\infty}^{\infty}dr\,p_{R}(r)r^{n},$ (41)
where the distribution $p_{R}(r)$ reads
$p_{R}(r)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum\limits_{i=1}^{N}\delta(r-R_{1,i}).$ (42)
The moments of $p_{R}(r)$ characterize the fluctuations of the eigenvector components, allowing to study localization phenomena. In reference Metz and Neri (2021), the first moments of $p_{R}(r)$ have been analytically computed for sparse directed networks, unveiling a localization transition as a function of the network parameters.
IV.2 Phase diagram
Before discussing how the epidemic threshold depends on the network parameters, we compare the solutions of Eq. (11), valid for $N\rightarrow\infty$, with results obtained from the fixed-point Eq. (8) on finite-sized networks. Figure 1 shows numerical results for the prevalence $\langle\rho\rangle$ and the full distribution $\mathcal{P}(\rho)$. In both cases, the agreement between the solutions of Eq. (11) and those obtained from Eq. (8) is excellent. The distribution $\mathcal{P}(\rho)$ features a Dirac-$\delta$ at $\rho=0$, reflecting a finite fraction of nodes with zero indegree. While both approaches rely on numerical computations, an important advantage of Eq. (11) over Eq. (8) is that the former does not require the use of sophisticated algorithms to sample networks from the configuration model, since it depends on the network structure only through $p_{{\rm in},k}$ and $P_{\lambda}(x)$.
Figure 1: Comparison between the solutions of Eq. (11) (solid lines) with the fixed-point solutions of Eq. (8) (symbols) for directed networks with a Poisson degree distribution and a $\Gamma$-distribution of infection rates or coupling strengths. The fixed-point solutions are derived from an ensemble of 5050 networks with $N=10^{5}$ nodes, while the population dynamics results are obtained from 55 independent runs of the algorithm with $M=10^{6}$ stochastic variables. (a) The prevalence $\langle\rho\rangle$ as a function of the mean infection rate $\lambda$ for standard deviation $\sigma=0.2$ of the infection rates. (b) and (c): distribution $\mathcal{P}(\rho)$ of the infection probabilities for $\lambda=1/2$, $\sigma=0.2$, and two different $c$.
In figure 2, we present the phase diagram of the SIS model on directed networks in terms of $(\sigma,c)$. The model exhibits an absorbing phase, where $\mathcal{P}(\rho)=\delta(\rho)$, and an endemic phase, characterized by a stationary distribution $\mathcal{P}(\rho)$ with nonzero prevalence $\langle\rho\rangle$. The average degree $c=\lambda^{-1}$ above which the epidemic spreads to a finite fraction of the population is determined solely by the mean infection rate and is independent of other network properties. If $(\sigma,c)$ lies above the dashed line in the phase diagram, the spectral gap $|\Lambda_{1}-\Lambda_{2}|$ is finite; otherwise, $|\Lambda_{1}-\Lambda_{2}|=0$. For $\sigma<\sigma_{*}$, the transition line that delimits the absorbing phase follows from the linear stability analysis of the disease-free fixed-point. For $\sigma>\sigma_{*}$, the transition line is obtained by numerically solving Eq. (11) and monitoring the prevalence $\langle\rho\rangle$. The colour scale in figure 2 quantifies the inverse participation ratio of the infection probabilities, which will be discussed in the next subsection.
The linear stability analysis identifies the leading eigenvector $\bm{R}_{1}$ as responsible for destabilizing the absorbing phase. By the Perron-Frobenius theorem Horn and Johnson (2012), the components of $\{R_{1,i}\}_{i=1}^{N}$ are non-negative. Above the dashed line in figure 2, $\bm{R}_{1}$ is associated with an outlier eigenvalue, characterized by $\langle R_{1}\rangle>0$ Metz and Neri (2021). This “ferromagnetic” mode triggers the onset of the endemic phase. Below the dashed line, the leading eigenvalue lies at the boundary of the continuous spectrum, where $\langle R_{1}\rangle=0$ Metz and Neri (2021). Combined with the constraint $R_{1,i}\geq 0$, this suggests that $R_{1,i}=0$ with probability one in the limit $N\rightarrow\infty$. This mode is unable to destabilize the absorbing state $\mathcal{P}(\rho)=\delta(\rho)$, which explains the absence of a transition to the endemic phase for $c<\lambda^{-1}$, in contrast with the prediction of Eq. (40). Hence, the phase transition for $\sigma>\sigma_{*}$ in figure 2 is not governed by the leading eigenpair of the contact network.
Figure 2: Phase diagram of the SIS model on directed networks in terms of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates or coupling strengths (the mean infection rate is $\lambda=1/2$). The indegrees follow a Poisson distribution, while the infection rates follow a $\Gamma$-distribution. The model exhibits an endemic phase ($\langle\rho\rangle>0$) and an absorbing phase ($\langle\rho\rangle=0$). The standard deviations at the dots are, respectively, $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$. For $\sigma>\sigma_{*}$, the critical line is obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables. The colour scale shows the inverse participation, Eq. (45), which quantifies the spatial fluctuations of the infection probabilities. The fourth moment $\mu_{4}$ of the infection rates is defined in Eq. (49).
Figure 3: Prevalence $\langle\rho\rangle$ as a function of the mean degree $c$ for different distributions of infection rates and indegrees in the regime $\sigma<\sigma_{*}$ (see the main text). The infection rates or coupling strengths have mean $\lambda=1/2$ and standard deviation $\sigma=0.2$. The results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{5}$ stochastic variables. The inset shows the prevalence near $c=\lambda^{-1}$ in logarithmic scale. The colours in the inset correspond to the same distributions as in the main panel.
Figure 4: Prevalence $\langle\rho\rangle$ as a function of the standard deviation $\sigma$ of the infection rates for different distributions of indegrees and infection rates in the regime $\sigma>\sigma_{*}$ (see the main text). The average indegree is $c\simeq 2.7$, and the infection rates have mean $\lambda=1/2$. For power-law distributed indegrees, the smallest indegree is $k_{\rm min}=2$. The results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables.
We now discuss the role of network heterogeneities on the prevalence $\langle\rho\rangle$ and the epidemic threshold. Figure 3 shows $\langle\rho\rangle$ as a function of $c$ for different distributions $p_{{\rm in},k}$ and $P_{\lambda}(x=\lambda_{ij})$ in the regime $\sigma<\sigma_{*}$. The prevalence vanishes as $\langle\rho\rangle\simeq c-\lambda^{-1}$ for $0<c-\lambda^{-1}\ll 1$, consistent with the mean-field critical exponent of directed percolation Ódor (2004). The epidemic threshold as well as the critical behaviour of $\langle\rho\rangle$ are both independent of the distributions $p_{{\rm in},k}$ and $P_{\lambda}(x)$, confirming the universality of the transition at $c=\lambda^{-1}$.
In the regime $\sigma>\sigma_{*}$, the stationary behaviour becomes highly sensitive to the shape of the distribution $P_{\lambda}(x)$ of infection rates. As shown in figure 4, when $\lambda_{ij}$ follows a $\Gamma$-distribution, the prevalence drops to zero at sufficiently large $\sigma$, due to the large fraction of small infection rates caused by the divergence of $P_{\lambda,{\rm g}}(x=0)$ (see Eq. (4)). In contrast, when $\lambda_{ij}$ follows a Pareto distribution, in which the smallest infection rate is $x_{0}>0$, the endemic state persists even for strong fluctuations of $\lambda_{ij}$, and $\langle\rho\rangle$ saturates at a finite value. Taken together, figures 3 and 4 show that the epidemic threshold is independent of the indegree distribution $p_{{\rm in},k}$.
IV.3 Localization of epidemic spreading
In this section we present results for higher moments of $\mathcal{P}(\rho)$, which characterize the spatial fluctuations of the infection probabilities and the localization of the epidemics. We also discuss the connection between the moments of $\mathcal{P}(\rho)$ and those of the distribution of the leading eigenvector $\bm{R}_{1}$.
The inverse participation ratio (IPR) is a standard probe of spatial localization in disordered systems Mirlin (2000); Goltsev et al. (2012); Metz and Neri (2021). Following Metz and Neri (2021), the IPR of the fixed-point vector $\bm{\rho}=(\rho_{1},\dots,\rho_{N})^{T}$ is defined as follows
$\mathcal{I}_{N}^{(4)}(\bm{\rho})=\dfrac{N\sum_{i=1}^{N}\rho_{i}^{4}}{\left(\sum_{i=1}^{N}\rho_{i}^{2}\right)^{2}},$ (43)
while the dimensionless second moment reads
$\mathcal{I}_{N}^{(2)}(\bm{\rho})=\dfrac{N\sum_{i=1}^{N}\rho_{i}^{2}}{\left(\sum_{i=1}^{N}\rho_{i}\right)^{2}}.$ (44)
If the network has a finite number of nodes with nonzero infection probabilities, $\mathcal{I}_{N}^{(4)}(\bm{\rho})$ and $\mathcal{I}_{N}^{(2)}(\bm{\rho})$ both scale linearly with $N$ and the vector $\bm{\rho}$ is localized. If the infection probabilities are nonzero on an extensive number of nodes, $\bm{\rho}$ is delocalized or extended, implying that $\mathcal{I}_{N}^{(4)}(\bm{\rho})$ and $\mathcal{I}_{N}^{(2)}(\bm{\rho})$ are of order $\mathcal{O}(N^{0})$. Here, we study the behaviour of these quantities strictly in the limit $N\rightarrow\infty$ by numerically solving Eq. (11) for $\mathcal{P}(\rho)$.
In the endemic phase, the disease infects a finite fraction of individuals and the above parameters converge to
$\mathcal{I}^{(4)}(\rho)=\lim_{N\rightarrow\infty}\mathcal{I}_{N}^{(4)}(\bm{\rho})=\dfrac{\langle\rho^{4}\rangle}{\langle\rho^{2}\rangle^{2}}$ (45)
and
$\mathcal{I}^{(2)}(\rho)=\lim_{N\rightarrow\infty}\mathcal{I}_{N}^{(2)}(\bm{\rho})=\dfrac{\langle\rho^{2}\rangle}{\langle\rho\rangle^{2}}.$ (46)
In figure 2, we quantify the fluctuations of the infection probabilities in the endemic phase by displaying $\mathcal{I}^{(4)}(\rho)$ in a colour scale. While the IPR can increase by several orders of magnitude near the phase transition, $\mathcal{I}^{(4)}(\rho)$ remains finite, indicating that the state vector $\bm{\rho}$ is delocalized in the endemic phase.
Nevertheless, the prevalence $\langle\rho\rangle$ vanishes continuously as we approach the critical line, suggesting that the disease may become localized near the epidemic threshold. To examine this scenario in more detail, we focus on the regime $\sigma<\sigma_{*}$, where the epidemic threshold is known analytically. Moreover, for $0<c-\lambda^{-1}\ll 1$ ($\sigma<\sigma_{*}$), $\bm{\rho}$ is governed by the leading eigenvector $\bm{R}_{1}$, responsible for destabilizing the absorbing phase. The moments of $\bm{R}_{1}$ have been analytically computed for directed complex networks in the limit $N\rightarrow\infty$ Metz and Neri (2021).
Figure 5: Dimensionless second moment $\mathcal{I}^{(2)}(\rho)$ [Eq. (46)] and inverse participation ratio $\mathcal{I}^{(4)}(\rho)$ [Eq. (45)] as functions of the standard deviation $\sigma$ of the infection rates for different mean degrees $c$ close to the epidemic threshold $c=\lambda^{-1}$. These results are for directed networks with Poisson indegrees and a $\Gamma$-distribution of infection rates or coupling strengths with mean $\lambda=1/2$. Symbols represent numerical results obtained from Eq. (11) using the population dynamics algorithm with $M=10^{6}$ (vertical bars indicate the standard deviation of the mean computed over 1010 independent runs). Solid lines are analytic predictions derived from the moments of the leading eigenvector [Eqs. (47) and (48)].
For $\sigma<\sigma_{*}$, the leading eigenvalue is an outlier and the ratio $\mathcal{I}^{(2)}(R_{1})$ of the first two moments of $\bm{R}_{1}$ for a Poisson indegree distribution fulfills Metz and Pérez Castillo (2016); Lucas Metz et al. (2019); Neri and Metz (2020)
$\mathcal{I}^{(2)}(R_{1})=\dfrac{c}{c-1-\sigma^{2}/\lambda^{2}},$ (47)
while the IPR of $\bm{R}_{1}$ is given by
$\mathcal{I}^{(4)}(R_{1})=\dfrac{c^{3}\left(3\mu_{3}+4c\mu_{2}+c^{2}\right)-2c\mu_{2}^{2}\left(c^{2}+3\mu_{3}\right)}{\left(c^{3}-\mu_{4}\right)\left(c^{2}-\mu_{3}\right)},$ (48)
where we defined the dimensionless moments
$\mu_{n}=\lambda^{-n}\int_{0}^{\infty}dxP_{\lambda}(x)x^{n}$ (49)
of the infection rates or coupling strengths. Depending on the network parameters, Eqs. (47) and (48) diverge due to vanishing denominators Metz and Neri (2021). In particular, $\mathcal{I}^{(4)}(R_{1})$ diverges at $c=\mu_{4}^{\frac{1}{3}}$. This naturally raises the question of whether $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ exhibit a similar behaviour near $c=\lambda^{-1}$.
In figure 5, we show $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ as functions of $\sigma$ near the epidemic threshold. As a comparison, this figure also displays Eqs. (47) and (48) at $c=\lambda^{-1}$, demonstrating that $\mathcal{I}^{(2)}(R_{1})$ and $\mathcal{I}^{(4)}(R_{1})$ diverge at $\sigma_{*}=\sqrt{\lambda(1-\lambda)}$ and $\sigma_{4}\simeq 0.36$, respectively, for the $\Gamma$-distribution of infection rates. The results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ are overall consistent with the analytic predictions obtained from the moments of $\bm{R}_{1}$. The discrepancies observed in figure 5 stem from the slow convergence of the population dynamics algorithm near the thresholds $\sigma_{*}$ and $\sigma_{4}$, which makes it difficult to accurately determine the distribution $\mathcal{P}(\rho)$ from the solutions of Eq. (11), as shown in appendix A. In addition, we note that $\mathcal{I}^{(4)}(\rho)$ remains bounded for $\sigma_{4}<\sigma<\sigma_{*}$, a consequence of the finite number of stochastic variables used to discretize $\mathcal{P}(\rho)$ in the population dynamics method.
Figure 6: Probability density $\mathcal{P}_{Y}(y)$ of $Y_{i}=-\ln\rho_{i}$ ($i=1,\dots,N$) near the epidemic threshold ($c-\lambda^{-1}=10^{-4}$). The directed network is characterized by a Poisson indegree distribution and a $\Gamma$-distribution of infection rates with mean $\lambda=1/2$ and varying standard deviation $\sigma$. The results are obtained from the numerical solutions of Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables. $\mathcal{P}_{Y}(y)$ exhibits an exponential decay for large $y$.
While the moments of the infection probabilities vanish as we approach the critical line, the ratios $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ diverge for sufficiently large $\sigma$, suggesting that $\bm{\rho}$ may become localized slightly above the epidemic threshold. This divergence arises from a large fraction of nodes with $\rho_{i}\simeq 0$ and a small fraction of nodes with comparatively larger $\rho_{i}$. To gain further insight into the fluctuations of $\rho_{i}$, we show in figure 6 the probability density $\mathcal{P}_{Y}(y)$ of $Y_{i}=-\ln\rho_{i}\in[0,\infty)$ near the epidemic threshold. The exponential decay $\mathcal{P}_{Y}(y)\propto e^{-Ay}$ for large $y$ implies that $\mathcal{P}(\rho)\propto\rho^{A-1}$ for $0<\rho\ll 1$, where the exponent $A$ depends on the network parameters. The power-law behaviour of $\mathcal{P}(\rho)$ near $\rho=0$ reflects strong fluctuations of $\rho_{i}$ spanning several orders of magnitude, yet confined to very small values of $\rho_{i}$.
IV.4 Stochastic simulations
In this section, we compare our theoretical results in the limit $N\rightarrow\infty$, obtained within the QMF approximation, with numerical results from stochastic simulations of the SIS model Kiss et al. (2017); Cota and Ferreira (2017); Wu et al. (2019) on directed networks. The simulations are performed using the well-known Gillespie algorithm Kiss et al. (2017).
In the stochastic simulations of the SIS model, the dynamical state of each node $i$ is represented by a binary variable $s_{i}\in\{0,1\}$ ($i=1,\dots,N$). If node $i$ is infected, then $s_{i}=1$; otherwise, $s_{i}=0$. The uniform recovery rate $R_{i}$ of node $i$ is given by $R_{i}=s_{i}$, while the infection rate of vertex $i$ reads
$I_{i}=(1-s_{i})\sum_{j\in\partial_{i}^{\rm in}}\lambda_{ij}s_{j},$ (50)
where $\partial_{i}^{\rm in}$ denotes the in-neighbourhood of $i$. The coefficient $\lambda_{ij}>0$ is the directed coupling strength from node $j$ to $i$, representing the rate at which $j$ infects $i$. The variables $\{\lambda_{ij}\}_{i,j=1,\dots,N}$ are sampled from a distribution $P_{\lambda}(x)$ which, together with the degree distribution $p_{k,\ell}$, defines the network ensemble (see section II). Thus, infected nodes recover at a unit and constant rate, while susceptible nodes may become infected through contact with their in-neighbourhood.
The time increment between two consecutive updates in the simulation is sampled from an exponential distribution with rate equal to the total rate of change, $\sum_{i=1}^{N}(R_{i}+I_{i})$. To perform a single update in the system, we first select the type of event according to the total recovery and infection rates. If the selected event is a recovery, an infected node is chosen at random and its state is updated to susceptible. If the event is an infection, a susceptible node becomes infected with a probability proportional to its infection rate. To select which susceptible node becomes infected, we employ a rejection sampling method, i.e., we choose a node at random and accept the choice with probability $I_{i}/\sum_{j=1}^{N}I_{j}$. The prevalence at a given time is computed as
$\langle s\rangle=\frac{1}{N}\sum_{i=1}^{N}s_{i}.$ (51)
The stochastic dynamics of the SIS model on a finite network always converges to the absorbing state for sufficiently long times Sander et al. (2016); Cota and Ferreira (2017), which leads to computational difficulties in determining the epidemic threshold. To circumvent this issue, we employ the quasistationary method Sander et al. (2016), where the prevalence at time $t$ is computed by averaging Eq. (51) over the subset of realizations that have not reached the absorbing state.
Figure 7 compares theoretical results obtained from the solutions of Eq. (11) with stochastic simulations for different system sizes $N$. Finite size effects are small in the simulations, and the theoretical predictions for the prevalence, based on the QMF approximation, are in qualitative agreement with the simulation results. Since the QMF equations neglect pairwise dynamical correlations, the prevalence computed from Eq. (11) is systematically larger than the simulation results. Accordingly, the theoretical epidemic threshold is consistently smaller than the one observed in simulations. Nevertheless, figure 7 shows that the agreement between simulations and the QMF approximation improves as either $c$ increases or $\sigma$ decreases, i.e., in regimes where fluctuations in the contact network structure become less pronounced.
Figure 7: Comparison between theoretical results in the limit $N\rightarrow\infty$ (dashed lines) and stochastic simulation results (symbols) for the stationary prevalence of the SIS model on directed networks. The indegrees and outdegrees are independently drawn from a Poisson distribution with mean $c$, while the infection rates or coupling strengths $\lambda_{ij}$ follow a $\Gamma$-distribution with mean $\lambda=1/2$ and standard deviation $\sigma$. The theoretical results are obtained by solving Eq. (11) using the population dynamics algorithm with $M=10^{6}$ stochastic variables, while the simulation results are computed by averaging the prevalence over 1010 independent realizations after a long simulation time. The simulations are performed using the Gillespie algorithm (see the main text). Left panel (a): $\sigma=0.2$. Right panel (b): $c=5$.
Figure 8: Stochastic simulation results (symbols) for the prevalence of the SIS model on directed networks with two different degree distributions and a $\Gamma$-distribution of infection rates with mean $\lambda=1/2$ and standard deviation $\sigma$. The simulations are performed using the Gillespie algorithm (see the main text), and the results are obtained by averaging the prevalence over 110110 independent realizations of size $N=10^{4}$ after a long simulation time. The solid lines are a guide to the eye. Left panel (a): $\sigma=0$. Right panel (b): $c=3$.
In figure 8, we present simulation results for the prevalence across the epidemic threshold for $N=10^{4}$ and two different degree distributions. These results suggest that the epidemic threshold is indeed independent of the degree distribution, in qualitative agreement with the predictions derived from Eq. (11) (see figures 3 and 4). Thus, even though our theoretical findings are based on the QMF approximation, Eqs. (2), this simplified description appears to capture some essential aspects of the phase diagram.
V Final remarks
In this work, we have determined the phase diagram of the SIS model on directed complex networks within the quenched mean-field approximation. By combining random-matrix results with an analytic approach for the distribution of stationary infection probabilities, we have computed the epidemic threshold as a function of the mean degree $c$ and the standard deviation $\sigma$ of the infection rates or coupling strengths defining the contact network. Our results show that the SIS model exhibits a transition between the absorbing and endemic phases provided $c\geq\lambda^{-1}$, where $\lambda$ is the average infection rate.
Remarkably, the critical line is independent of the degree distribution but it depends strongly on the distribution of infection rates. When $\sigma<\sigma_{*}=\sqrt{\lambda(1-\lambda)}$, both the epidemic threshold and the critical behaviour are governed by the leading eigenpair of the contact network Metz and Neri (2021) and are independent of the infection-rate distribution. In contrast, for $\sigma>\sigma_{*}$, the fluctuations of infection rates have a pronounced effect on the critical line. While the SIS model undergoes an absorbing phase transition for a $\Gamma$-distribution of infection rates as $\sigma$ increases, it remains in the endemic phase for a Pareto distribution. This striking difference is explained by the large fraction of near-zero infection rates generated by the $\Gamma$-distribution, which suppresses the endemic state for large $\sigma$. Together, these results provide a systematic characterization of how network heterogeneity shapes the phase diagram of the SIS model on directed networks in the quenched mean-field approximation.
We have also examined the emergence of disease localization right above the epidemic threshold Goltsev et al. (2012). Focusing on the regime where the phase transition is governed by the leading eigenpair of the contact network, we have shown that the inverse participation ratio (IPR) of the fixed-point infection probabilities diverges near the threshold for sufficiently large $\sigma$, suggesting that the disease becomes localized on a vanishing fraction of nodes. These findings are consistent with analytic predictions based on the IPR of the leading eigenvector Metz and Neri (2021). We have also computed the full probability density of infection probabilities near the critical line and found that it exhibits a large fraction of near-zero values, with strong fluctuations spanning several orders of magnitude.
We have compared our theoretical results for $N\rightarrow\infty$ with stochastic simulations performed using the Gillespie algorithm Kiss et al. (2017); Cota and Ferreira (2017). The epidemic threshold computed from the solutions of Eq. (11) consistently lies below the value obtained from simulations. This quantitative discrepancy is not unexpected, since Eq. (11) is derived from the quenched mean-field (QMF) approximation, which neglects pairwise dynamical correlations Mata and Ferreira (2013). Nevertheless, the simulation results suggest that the exact epidemic threshold is independent of the degree distribution, in agreement with our theoretical predictions. Overall, this comparison with stochastic simulations indicates that the QMF approximation captures the main qualitative features of the phase diagram.
We remark that our results are strictly valid in the limit $N\rightarrow\infty$, as they rely on the numerical solutions of Eq. (11). Consequently, our approach does not address how the infection probabilities or the number of potentially infected nodes in the QMF theory scale with the system size, which could in principle clarify, for instance, why the dimensionless second moment remains finite while the IPR diverges. A detailed analysis of this issue and its relation with the epidemic survival time are interesting directions for future work. Furthermore, it would be interesting to unveil the mechanism sustaining the endemic state in directed graphs and compute the finite-size epidemic threshold by developing, for instance, a directed version of the scaling theory in Castellano and Pastor-Satorras (2020).
The SIS model on synthetic networks provides a theoretical framework to qualitatively investigate the effect of the heterogeneous structure of contact networks on epidemic spreading. However, we emphasize that, from a practical standpoint, this setting is highly idealized and insufficient for making quantitative predictions about real-world spreading phenomena. From a methodological perspective, we have introduced a general analytic framework to obtain an equation for the distribution of fixed-point states in the limit $N\rightarrow\infty$. The method is not restricted to the SIS model and can be extended to study the non-equilibrium fixed-points of other dynamical systems on directed networks, including the Kuramoto model Acebrón et al. (2005) and firing-rate models of neural networks Sompolinsky et al. (1988). Moreover, since the approach is based on the cavity method, it can also incorporate other important network features, such as short loops and degree-degree correlations, opening promising avenues of future research.
Acknowledgements.
V. B. M. acknowledges fellowships from CNPq/Brazil and BIC-UFRGS. F. L. M. acknowledges support from CNPq (Grant No 402487/2023-0), FAPERJ (Grant No 204.646/2024), and from ICTP through the Associates Program (2023-2028)
Appendix A Convergence of the population dynamics algorithm
In this appendix, we present numerical results for the influence of both the population size $M$ and the number of iterations $N_{\rm iter}$ in the population dynamics algorithm near the transition point $\sigma_{4}\simeq 0.36$ of figure 5. Figure 9 shows the difference between the analytic results for $N\rightarrow\infty$, Eqs. (47) and (48), and the numerical results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$ obtained from the solutions of Eq. (11). For relatively small number of iterations, there is a significant discrepancy between analytic and numerical results, which is not evident in figure 5 due to the logarithmic scale of the vertical axis. However, the data in figure 9 indicate that, despite the fluctuations, the numerical results consistently converge to the theoretical predictions as we increase both $M$ and $N_{\rm iter}$, giving further support to the analytic curves in figure 5. This critical slowing down of the population dynamics method close to $\sigma_{4}$ makes accurate calculations of higher moments of $\mathcal{P}(\rho)$ computationally very expensive.
Figure 9: Dimensionless second ($\mathcal{I}^{(2)}(\rho)$) and fourth ($\mathcal{I}^{(4)}(\rho)$) moments of the infection probabilities. The panels show the difference between the analytic results, $\mathcal{I}^{(2)}(R_{1})$ and $\mathcal{I}^{(4)}(R_{1})$ (see Eqs. (47) and (48)), and the numerical results for $\mathcal{I}^{(2)}(\rho)$ and $\mathcal{I}^{(4)}(\rho)$, obtained from the solutions of Eq. (11) using the population dynamics algorithm with population size $M$ and total number $N_{\rm iter}$ of iterations (vertical bars indicate the standard deviation of the mean values computed over different samples). These results are for parameters near the critical point in figure 5: Poisson indegrees, $\Gamma$-distribution of infection rates with mean $\lambda=1/2$ and standard deviation $\sigma=0.3$, and $c-\lambda^{-1}=10^{-4}$.